### v2

# Bipartite Fluctuations of Critical Fermi Surfaces

Abstract
Fluctuations of conserved quantities within a subsystem are non-local observables that provide unique insights into quantum many-body systems. In this paper, we explore the behaviors of bipartite charge (and spin) fluctuations across interaction-driven “metal-insulator transitions” out of Landau Fermi liquids. The “charge insulators” include a class of non-Fermi-liquid states of fractionalized degrees of freedom, such as compressible composite Fermi liquids (for spinless electrons) and incompressible spin-liquid Mott insulators (for spin-$1/2$ electrons). We find that charge fluctuations $\mathcal{F}$ exhibit distinct leading-order scalings across the phase transition: $\mathcal{F}\sim L\log(L)$ in Landau Fermi liquids and $\mathcal{F}\sim L$ in charge insulators, where $L$ is the linear size of the subsystem under bipartition. In the case of composite Fermi liquids (under certain conditions), we also identify a universal constant term $-\textrm{f}(\theta)|\sigma_{xy}|/(2\pi)$ arising when the subsystem geometry incorporates a sharp corner. Here, $\textrm{f}(\theta)$ represents a function of the corner angle, and $\sigma_{xy}$ denotes the Hall conductivity. At the critical point of each example, provided the transition is continuous, we find that the leading-order scaling $\mathcal{F}\sim L$ is accompanied by a subleading universal corner contribution $-\log(L)\textrm{f}(\theta)C_{\rho}/2$ with the same angle dependence $\textrm{f}(\theta)$. The universal coefficient $C_{\rho}$ is linked to the predicted universal longitudinal (and Hall) resistivity jump $\Delta\rho_{xx}$ (and $\Delta\rho_{xy}$) at the critical point.
I Introduction
Measuring bipartite fluctuations of local observables provides a powerful approach to characterizing entanglement properties, symmetries, and correlations in various quantum phases and phase transitions. Although entanglement has become a fundamental organizing principle for the study of quantum matter [1, 2], experimental measurement of many-body entanglement remains a challenge. Significant progress (see, e.g., Ref. [3, 4, 5, 6, 7, 8, 9, 10, 11, 12]) has been made toward establishing a direct relationship between entanglement entropy and the fluctuations of globally conserved quantities, such as particle number and magnetization. This relationship has been demonstrated to be feasible in special cases, such as one-dimensional Luttinger liquids [6, 8] and free fermions in higher dimensions [9, 8, 4]. However, the connection between the two in strongly correlated systems beyond these examples is still poorly understood. The two quantities are observed to differ from each other in continuous symmetry-breaking phases, where the entanglement entropy still follows a boundary-law scaling while the bipartite fluctuations exhibit a multiplicative logarithmic enhancement to the boundary law. Building on this observation, Ref. [13] proposed the use of the distinct scalings of bipartite fluctuations to detect conventional Landau phase transitions.
Despite decades of study, understanding unconventional quantum phases and phase transitions that are beyond Landau’s symmetry-breaking paradigm remains a central problem in condensed matter physics. In recent years, conceptual breakthroughs [14, 15] have been achieved by employing extended operators to define generalized symmetries and to characterize phases and phase transitions (see Ref. [16, 17] and references therein). For instance, abelian topological orders can be understood in terms of the condensation of 1-dimensional objects, leading to the spontaneous breaking of discrete $1$-form symmetries. Similarly, the Coulomb phase of electrodynamics can be interpreted as a spontaneous symmetry-breaking phase of $\textrm{U}(1)$ $1$-form symmetry, where gapless photons serve as Goldstone modes. This helps establish a generalized Landau symmetry paradigm. In this case, the criterion for determining whether the phase is symmetric or not is given by the scaling of disorder operators (or Wilson loops under duality). In this context, the concept of bipartite fluctuations is again closely related, which can be identified as the $\textrm{U}(1)$ disorder operator under the small-angle limit [18, 19].
While extended operators play an important role in the conceptual understanding of phases of matter, recent studies have also shed light on intriguing quantitative aspects of disorder operators (and related bipartite fluctuations), particularly their scaling behaviors at conformally invariant quantum critical points [20, 18, 19, 21, 22, 23, 24, 25, 26]. For instance, when considering the configuration of the subsystem as depicted in Fig. 1, bipartite fluctuations exhibit a universal corner contribution with logarithmic scaling. This contribution has a universal angle dependence and is directly proportional to the current central charge $C_{J}$ of the conformal field theory (CFT) [20, 18, 19, 21]. Given the sometimes uncertain fate of proposed lattice realizations for exotic quantum critical points, numerical simulations, such as quantum Monte Carlo simulations, can assist in identifying whether these lattice models correspond to unitary CFTs by examining the sign of the universal corner contribution [22, 23, 24, 25, 26]. Moreover, Ref. [21] showed that the corner term of bipartite fluctuations in isotropic integer quantum Hall (QH) and Laughlin states is determined by the DC Hall conductivity.
Understanding quantum phase transitions in metals is significantly more challenging due to the abundance of low-energy excitations near the electronic Fermi surface. Even for conventional phase transitions associated with some form of broken symmetry in metals, the standard Hertz-Millis-Moriya framework [27, 28, 29] encounters serious difficulties in two spatial dimensions. Despite numerous attempts made in recent years [30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47], many aspects of the low-energy physics remain difficult to describe theoretically. A scenario that is both technically and conceptually more challenging is a continuous transition from an ordinary metal to an exotic gapless phase with abundant fractionalized excitations forming a Fermi surface. One such example is the proposed continuous Mott transition at half-filling [48, 49, 50, 51] from a Landau Fermi liquid (FL) to a gapless Mott insulator (MI) with a spinon Fermi surface. This is potentially realized by the recent experimental observation [52] of a continuous bandwidth-tuned transition from a metal to a paramagnetic Mott insulator in the transition metal dichalcogenide (TMD) Moiré heterobilayer $\textrm{MoTe}_{2}/\textrm{WSe}_{2}$. However, the observed critical resistivity is anomalously large [52], exceeding the predictions of the original theory, at least in the clean limit [^1] See Ref. [53] for a discussion about the effects of disorder.. To address this discrepancy, Ref. [54] has proposed a modified theory that explains the large critical resistivity by charge fractionalization.
Another notable example that has appeared in literature is the bandwidth-tuned transition from a Landau Fermi liquid (FL) to a composite Fermi liquid (CFL) describing the half-filled Landau level. The theoretical possibility of a continuous transition was originally proposed by Ref. [55] and later refined by Ref. [56] targeting TMD Moiré materials [^2] The existence of CFLs at 1/2 and 1/4 fillings in twisted $\textrm{MoTe}_{2}$ is supported by Ref. [57, 58].. A general criterion [59, 60, 61, 62] distinguishing charge insulators from metals is the presence of a finite localization tensor, i.e., finite polarization fluctuations. Accordingly, as will be illustrated below, the CFL may be viewed as a “charge insulator,” and the CFL-FL transition can be interpreted as a “metal-insulator transition” in a broad sense.
All these “metal-insulator transitions” share a common theoretical framework, which we will review in Sec. III, where electrons are fractionalized into fermionic and bosonic partons. The fermion sector forms a stable Fermi-surface state, while the boson sector contains a single relevant operator that drives the transition from a superfluid to a gapped state. Crucial insight into the possible transitions comes from the filling constraints of the bosons under translation symmetry and the Lieb-Schultz-Mattis theorem [54, 56]. Without better terminology, we refer to this type of quantum critical points as critical Fermi surfaces, following Refs. [50, 51].
Figure 1: The real-space subsystem $\mathcal{A}$ with only one corner angle $0<\theta<\pi$. The linear size of $\mathcal{A}$ is denoted by $L$.
In this work, we point out the universal features of critical Fermi surfaces from the perspective of bipartite charge fluctuations $\mathcal{F}$. Although the transitions go beyond any symmetry principles, including the generalized ones [16, 17], the two phases can be distinguished by the distinct leading-order scalings $\mathcal{F}\sim L$ and $\mathcal{F}\sim L\log(L)$, just like conventional symmetry-breaking transitions [13], where $L$ is the linear size of the subsystem under bipartition. Additionally, at the critical point, we find a subleading corner contribution with logarithmic scaling, reminiscent of the behavior seen in CFTs, despite the absence of conformal symmetry in such systems. The universal coefficient (denoted by $C_{\rho}$) of the logarithmic term can be directly linked to transport observables, such as the predicted longitudinal (and Hall) resistivity jump [51, 54, 56] at the critical point. Furthermore, based on an isotropic field theory [63], we predict a universal corner term in the non-Fermi-liquid phase of composite fermions governed by the DC Hall conductivity.
In contrast to conformally invariant quantum critical points, many powerful theoretical tools, such as the conformal bootstrap [64], are no longer applicable to strongly correlated metals. However, it is theoretically feasible in Monte Carlo simulations [22, 23, 24, 25, 26] to extract the subleading universal corner term of $\mathcal{F}$ using the method proposed in Ref. [65]. We anticipate that our findings, which establish a connection between the universal data $C_{\rho}$ of the critical points and charge fluctuations which are numerically and experimentally accessible, could aid future studies in identifying the existence and lattice realizations of these exotic quantum critical points in metals.
The remainder of the paper is organized as follows. In Sec. II, we provide relevant background and summarize our new findings. Sec. III introduces a unified theoretical framework for the continuous phase transitions discussed in this work. In Sec. IV, we explore spinless electrons, highlighting the intriguing behaviors of bipartite charge fluctuations within the CFL phase and at the CFL-FL transition. Sec. V shifts focus to spin-$1/2$ electrons, examining the universal behaviors of both charge and spin fluctuations across the Mott transitions. Finally, Sec. VI offers further discussion of our results and suggests promising directions for future research.
II Bipartite Fluctuations
In this section, we begin with a brief introduction to the preliminaries of bipartite fluctuations in various quantum states of matter, along with a general criterion for distinguishing charge insulators from metals. We then summarize our main findings on bipartite fluctuations across a class of bandwidth-tuned continuous “metal-insulator transitions.”
II.1 Preliminaries I
For any many-body system with a $\textrm{U}(1)$ global symmetry, the bipartite fluctuations can be defined as
$\mathcal{F}_{\mathcal{A}}=\left\langle\left(\int_{\mathcal{A}}J^{\tau}\right)-\left\langle\int_{\mathcal{A}}J^{\tau}\right\rangle\right\rangle^{2}.$ (1)
Here, $J^{\tau}$ represents the zeroth component of the conserved Noether current $J^{\mu}$ satisfying $\partial_{\mu}J^{\mu}=0$, and $\mathcal{A}$ is a spatial subregion within the total system. The expectation value is taken with respect to the ground state at zero temperature. By definition, the large-scale scaling behavior of $\mathcal{F}_{\mathcal{A}}$ is ultimately governed by the long-wavelength behavior of the equal-time density-density correlation $\Pi^{\tau\tau}=\langle J^{\tau}J^{\tau}\rangle$. In App. D, we undertake the task of evaluating $\mathcal{F}_{\mathcal{A}}$ assuming a generic power-law instantaneous charge correlation
$\Pi^{\tau\tau}(\tau\rightarrow 0,\bm{x})=\frac{C_{0}}{|\bm{x}|^{\alpha}}$ (2)
in two dimensions. The results are summarized in Tab. 1.
leading-order term $\epsilon$-independent term
$32<\alpha<$ 3 $L^{4-\alpha}$ $L^{4-\alpha}$
$\alpha=3$ $L\log(L)$ $L\log(L)$
$43<\alpha<$ 4 $L/\epsilon^{\alpha-3}$ $L^{4-\alpha}$
$\alpha=4$ $L/\epsilon$ $\log(L)$
$\alpha>4$ $L/\epsilon^{\alpha-3}$ 00
Table 1: The scaling behaviors of the bipartite fluctuations $\mathcal{F}_{\mathcal{A}}$ based on the density-density correlation Eq. (2). Here, $L$ represents the linear size of $\mathcal{A}$, and $\epsilon$ is a gauge-invariant UV cut-off introduced in the vortex theory (see App. D). There are some notable features. The leading-order term exhibits boundary-law scaling when $\alpha>3$. A universal subleading logarithmic term appears only when $\alpha=4$ and the geometry of $\mathcal{A}$ includes any sharp corners. For $\alpha>4$, the universal subleading term vanishes in the large-$L$ limit.
The simplest compressible state is given by the free fermi gas. Utilizing the free-fermion propagator, one can easily obtain $\Pi^{\tau\tau}(\tau\rightarrow 0,\bm{x})$ in real space or the static structure factor $\Pi^{\tau\tau}(\tau\rightarrow 0,\bm{k})$ in momentum space
$\Pi^{\tau\tau}(\tau\rightarrow 0,\bm{x})$ $=\frac{k_{F}}{4\pi^{3}}\frac{1}{|\bm{x}|^{3}},$
$\Pi^{\tau\tau}(\tau\rightarrow 0,\bm{k})$ $=-\frac{k_{F}|\bm{k}|}{2\pi^{2}},$ (3)
where $k_{F}$ denotes the fermi momentum. According to Tab. 1, we observe the scaling $\mathcal{F}\sim L\log(L)$. This is a well-known result and is closely related to entanglement entropy [66, 67, 68, 69, 9, 8, 4]. In App. A, we reproduce Eq. (3) based the $\textrm{LU}(1)$ anomaly of Fermi surface states [45], which holds even in the presence of strong forward scattering preserving the $\textrm{LU}(1)$ symmetry. The bonus is a geometric interpretation [70] of the static structure factor through the area difference when shifting the Fermi surface, as depicted in Fig. 2.
Figure 2: The geometric interpretation of the static structure factor $\Pi^{\tau\tau}(\tau\rightarrow 0,\bm{q})$ for Fermi surface states with LU(1) symmetry, where the shaded area represents its value.
Another simple compressible state is provided by the superfluid phase. Let us denote the $\textrm{U}(1)$ order parameter by $\phi$. The gapless Goldstone mode $\theta$ corresponds to the angle fluctuation of $\phi=\rho_{s}e^{\mathtt{i}\theta}$, where $\rho_{s}$ represents the stiffness. Since the power counting of the action $\int\textrm{d}^{3}x(\partial\theta)^{2}$ yields a scaling dimension of $\Delta[\partial\theta]=3/2$ for the $\textrm{U}(1)$ charge density, the corresponding equal-time correlation scales as $|\bm{x}|^{-3}$. Once again, one finds the scaling of bipartite fluctuations $\mathcal{F}\sim L\log(L)$.
An important class of incompressible states is given by unitary CFTs. The two-point function of the conserved spin-$1$ current has a rigid structure [64]
$\left\langle J^{\mu}(x)J^{\nu}(0)\right\rangle=\frac{C_{J}}{|x|^{4}}\left(\delta^{\mu\nu}-\frac{2x^{\mu}x^{\nu}}{|x|^{2}}\right),$ (4)
where the scaling dimension of $J^{\mu}$ is protected. The overall coefficient $C_{J}>0$, known as the current central charge, is a universal data of the CFT and is related to the longitudinal conductivity via $\sigma_{xx}=\frac{\pi^{2}}{2}C_{J}$. Considering the subsystem with a single corner, as shown in Fig. 1, one can verify the scaling behavior [18, 21, 19]
=#$\mathcal{F}=\#L-\frac{C_{J}}{2}\textrm{f}(\theta)\log(L)+\ldots.$ (5)
where ##\## is a non-universal number depending on the UV cut-off, and the universal angle dependence
$\textrm{f}(\theta)=1+(\pi-\theta)\cot(\theta).$ (6)
applies to any $(2+1)$-dimensional CFTs. This corner contribution has also been previously discussed in the context of non-interacting Dirac systems [20]. However, the question remains whether the universal behavior described in Eq. (5) could also exist at quantum critical points lacking conformal symmetry.
Insulators with a charge gap represent another standard type of incompressible states. Due to the exponential decay of the equal-time density-density correlation, the leading-order scaling obeys the boundary law. An intriguing observation from Ref. [21] is that certain gapped QH insulators, which possess continuous translational and rotational symmetries, exhibit a universal corner contribution when considering the subregion geometry shown in Fig. 1:
=#$\mathcal{F}=\#L-\frac{|\sigma_{xy}|}{2\pi}\textrm{f}(\theta)+\ldots.$ (7)
The angle dependence is once again given by Eq. (6), and $\sigma_{xy}=\nu/(2\pi)$ represents the DC Hall conductivity, where $\nu$ denotes the filling factor. The corner term for integer fillings $\nu\in\mathbb{Z}$ can be derived analytically, while the fractional QH insulator at $\nu=1/3$ has been verified using Monte Carlo simulations based on the Laughlin wavefunction [21]. It would be interesting to explore whether a similar corner term exists in compressible QH states at even-denominator fillings and to understand the conditions under which the formula in Eq. (7) holds.
II.2 Preliminaries II
FL MI CFL
Compressibility finite zero finite
Drude weight finite zero zero
Localization tensor divergent finite finite
Charge fluctuations $\mathcal{F}^{c}\sim L\log(L)$ $\mathcal{F}^{c}\sim L$ $\mathcal{F}^{c}\sim L$
Table 2: The distinct behaviors of various physical quantities across the Mott transition [50, 51, 54] from a Landau Fermi liquid (FL) to a Mott insulator (MI), and the transition [55, 56, 71] from a Landau FL to a composite Fermi liquid (CFL). In this work, we emphasize the different behaviors of the localization tensor (i.e., polarization fluctuations) and the leading-order scaling of charge fluctuations.
In this subsection, we clarify some terminology that will be useful in presenting our results. By definition, bipartite fluctuations depend on the inversion-symmetric part of the static structure factor
$\Pi_{+}^{\tau\tau}(\tau\rightarrow 0,\bm{k})=\frac{\Pi^{\tau\tau}(\tau\rightarrow 0,\bm{k})+\Pi^{\tau\tau}(\tau\rightarrow 0,-\bm{k})}{2}.$ (8)
Its long-wavelength expansion can be expressed as
$\Pi_{+}^{\tau\tau}(\tau\rightarrow 0,\bm{k})=\mathsf{G}^{ij}k_{i}k_{j}+\mathscr{O}(|\bm{k}|^{3}),$ (9)
where $\mathsf{G}^{ij}$ represents the real part of the localization tensor (i.e., polarization fluctuations) [59, 60, 61, 62]. Recently, $\mathsf{G}^{ij}$ has also been referred to as the “quantum weight” in Refs. [72, 73, 74]. In this work, we are interested in phase transitions from a Landau FL, where $\mathsf{G}^{ij}$ is divergent, to another phase where $\mathsf{G}^{ij}$ becomes finite. According to the general criterion in Ref. [59, 60, 61, 62], systems with a finite value of $\mathsf{G}^{ij}$ are classified as charge insulators.
Further physical insights into the localization tensor $\mathsf{G}^{ij}$ can be gained by considering the Souza-Wilkens-Martin (SWM) sum rule [60]
$\mathsf{G}^{ij}=\int_{0}^{+\infty}\frac{\textrm{d}\omega}{\pi}\frac{\textrm{Re}\sigma_{+}^{ij}(\omega)}{\omega},$ (10)
where $\sigma_{+}^{ij}=(\sigma^{ij}+\sigma^{ji})/2$ denotes the longitudinal conductivity. It establishes a fluctuation-dissipation theorem that connects the ground-state fluctuations of polarization to the optical conductivity. The linear behavior $\Pi^{\tau\tau}(\tau\rightarrow 0,\bm{k})\sim|\bm{k}|$ observed in Landau FLs (see also Eq. (3) for the free fermi gas) can be explained by the presence of a nonzero Drude weight $D^{ij}$ such that $\sigma_{+}^{ab}(\omega)\supset D^{ij}(\delta(\omega)+\frac{\mathtt{i}}{\pi\omega})$, leading to a linear divergence in $\mathsf{G}^{ij}\sim 1/|\bm{k}|$ due to the identical scaling of frequency $\omega$ and momentum $|\bm{k}|$. It is worth noting that a vanishing Drude weight does not necessarily imply a finite value of $\mathsf{G}^{ij}$. An example of this is provided by CFTs, which are incompressible, where the localization tensor diverges logarithmically as $\mathsf{G}^{ij}\sim\log(1/|\bm{k}|)$.
In Tab. 2, we summarize the behaviors of several physical quantities across the “metal-insulator transitions” where the “charge insulators” include both compressible and incompressible examples. Given that the primary focus of this work is on the universal field theories describing phase transitions, we assume continuous rotational symmetry for simplicity. This assumption renders the localization tensor isotropic, such that $\mathsf{G}^{ij}=\mathsf{G}\delta^{ij}$, where
$\mathsf{G}=\frac{\textrm{Tr}(\mathsf{G}^{ij})}{2}=\sqrt{\det(\mathsf{G}^{ij})}.$ (11)
Charge insulators with a general anisotropic $\mathsf{G}^{ij}$ will be addressed in an upcoming work [75].
II.3 Summary of Results
If the above-mentioned phase transitions are indeed continuous, the critical points are expected to be incompressible, with a vanishing Drude weight and intriguing universal behaviors in the localization tensor and charge fluctuations, which will be summarized below. The descriptions of these continuous phase transitions are unified within a theoretical framework that introduces vortex excitations in Landau FLs, as will be discussed in Sec. III. Different charge insulators can then be realized by placing vortices in different states.
Below, we summarize the key results in Sec. IV and Sec. V, where we consider examples of clean electron systems at half-filling, assuming disorder effects are weak.
Figure 3: (1) and (2) depict the predicted universal longitudinal resistivity jump $\Delta\rho_{xx}$ and Hall resistivity jump $\Delta\rho_{xy}$ at the transition between a Fermi liquid (FL) and a composite Fermi liquid (CFL) [56]; (3) illustrates the predicted universal longitudinal resistivity jump $\Delta\rho_{xx}$ at the transition between a Fermi liquid (FL) and a Mott insulator (MI) [51, 54]. In all of these phase transitions, the tuning parameter $g$ is related to the electron bandwidth, and the critical resistivity jumps are linked to the universal corner contribution of bipartite charge fluctuations, as described by Eq. (15).
  1. 1.
    Despite being compressible, CFLs follow the leading-order boundary-law scaling $\mathcal{F}\sim L$ of bipartite charge fluctuations, in contrast to the scaling $\mathcal{F}\sim L\log(L)$ seen in other compressible states like Landau FLs and superfluids. This difference arises because the instantaneous charge correlation from the gapless modes decays as $|\bm{x}|^{-5}$ (up to logarithmic corrections). Moreover, the gapped modes contribute a subleading corner term reminiscent of the behavior observed in gapped QH insulators [21]. Considering the subsystem geometry depicted in Fig. 1, the final result is
    =#$\mathcal{F}=\#L-\frac{\mathsf{G}}{\pi}\textrm{f}(\theta)+\ldots$ (12)
    where ##\## represents a non-universal constant, and $\mathsf{G}^{ij}=\mathsf{G}\delta^{ij}$ is the localization tensor introduced in Eq. (9) and Eq. (10). Based on the Halperin-Lee-Read (HLR) theory [63], we find the coefficient
    $\mathsf{G}=\frac{|\sigma_{xy}|}{2}$ (13)
    where $\sigma_{xy}=1/(4\pi)$ is the DC Hall conductivity of the half-filled Landau level. The angle dependence remains governed by the “super-universal” formula Eq. (6). Note that, in general, certain conditions must be satisfied for Eq. (13) to hold in charge insulators [75], as will be mentioned in Sec. IV.2.1.
  2. 2.
    At the critical point of the continuous CFL-FL transition, we find that the long-wavelength behavior of the static structure factor follows a CFT-like scaling, leading to a universal divergence of the localization tensor
    $\mathsf{G}=\frac{\pi C_{\rho}}{2}\log(\xi)+\ldots$ (14)
    where $\xi$ is the correlation length of the system. The coefficient $C_{\rho}$ should be understood as universal data associated with the critical Fermi surface [^3] We use a different notation $C_{\rho}$ for critical Fermi surfaces to distinguish it from $C_{J}$ used in CFTs.. An important phenomenology of this type of transition is the universal jumps in longitudinal resistivity $\Delta\rho_{xx}(\omega/T)$ and Hall resistivity $\Delta\rho_{xy}(\omega/T)$ at the critical point [56], where $\omega$ is frequency and $T$ is temperature, as depicted schematically in Fig. 3. In the limit $\omega/T\rightarrow\infty$, these jumps are related to the universal coefficient $C_{\rho}$ through
    $C_{\rho}=\frac{2}{\pi^{2}}\frac{\Delta\rho_{xx}}{(\Delta\rho_{xx})^{2}+(\Delta\rho_{xy})^{2}}.$ (15)
    Despite the absence of conformal symmetry in the system, the universal behavior in Eq. (14) results in the scaling of bipartite fluctuations similar to that seen in CFTs (see Eq. (5))
    =#$\mathcal{F}=\#L-\frac{C_{\rho}}{2}\textrm{f}(\theta)\log(L)+\ldots.$ (16)
  3. 3.
    The continuous Mott transition [51] can be detected by the distinct leading-order scalings of bipartite charge fluctuations: $\mathcal{F}^{c}\sim L\log(L)$ in the metallic phase and $\mathcal{F}^{c}\sim L$ in the insulating phase. At the non-conformal critical point, the localization tensor exhibits a universal divergence, as described in Eq. (14). The scaling of bipartite charge fluctuations $\mathcal{F}^{c}$ follows Eq. (16) with the coefficient $C_{\rho}$ given by the current central charge $C_{J}^{(\textrm{XY})}$ of the 3D XY universality class. The critical transport behavior is simpler due to the vanishing Hall response, and the relation in Eq. (15) to the critical resistivity jump still holds, with $\Delta\rho_{xy}=0$. The system also possesses a spin $\textrm{U}(1)$ symmetry, allowing for the definition of bipartite spin fluctuations $\mathcal{F}^{s}$. In both phases, the Fermi-surface states of spin excitations manifest as $\mathcal{F}^{s}\sim L\log(L)$. Richer physics can be revealed by generalizing the vortex theory in Sec. III to involve both charge and spin vortices. Further insights into the multicritical behaviors will be elaborated upon in Sec. V.2.
  4. 4.
    In the modified proposal [54] of the continuous Mott transition involving charge fractionalization, bipartite charge fluctuations $\mathcal{F}^{c}$ still exhibit distinct leading-order scalings of $\mathcal{F}^{c}\sim L\log(L)$ and $\mathcal{F}^{c}\sim L$ in the two phases. A key difference from the original theory [51] is observed in the universal behavior of $\mathcal{F}^{c}$ at the critical point, where it follows Eq. (16) but with a suppressed universal coefficient $C_{\rho}=2C_{J}^{(\textrm{XY})}/N^{2}$, where $C_{J}^{(\textrm{XY})}$ still represents the current central charge of the 3D XY fixed point. Here, $N$ is an even integer corresponding to the fractional electric charge $e/N$ carried by each charge carrier. Notably, in the time-reversal invariant version of Eq. (15) with $\Delta\rho_{xy}=0$, the critical resistivity jump $\Delta\rho_{xx}$ is enhanced by a factor of $N^{2}/2$, consistent with the experimentally observed large critical resistivity [52]. Another distinction from the original theory is seen in the bipartite spin fluctuations $\mathcal{F}^{s}$, which show boundary-law scaling $\mathcal{F}^{s}\sim L$ in the Mott insulator phase.
III Vortex Theory Framework
In this section, we present a general theoretical framework for describing a class of continuous quantum phase transitions out of Landau FLs, which will be utilized repeatedly in this paper.
One approach to quantum phase transitions is to incorporate vortex excitations into the Landau FL phase of electrons. If the vortex sector is not trivially gapped, the system is driven into a different phase. To formalize this idea, it is often convenient to begin with a parton decomposition of the spin-$1/2$ or spinless electron operator $c({\bm{x}})$, expressed as
$c({\bm{x}})=b({\bm{x}})f({\bm{x})}.$ (17)
where the assignment of the spin quantum number to the fermionic and bosonic partons $f$ and $b$ varies depending on specific examples, as summarized in Tab. 3. Notably, as will be discussed in Sec. VI, the same physics can be realized without introducing partons, by using nonlinear bosonization [47].
For illustrative purposes, let us focus on the case of spinless electrons. The generalization to spinful cases follows in a similar way. Due to the $\textrm{U}(1)$ gauge redundancy in Eq. (17), both $b$ and $f$ are coupled to a dynamical $\textrm{U}(1)$ gauge field $a_{\mu}$, carrying equal and opposite charges of ±$\pm 1$. When a background field $A_{\mu}$ is introduced for the global $\textrm{U}(1)$ symmetry, the effective Lagrangian can be schematically expressed as
$\mathcal{L}=\mathcal{L}[f,a+e_{f}A]+\mathcal{L}[b,-a+e_{b}A]+\ldots$ (18)
where $e_{f}+e_{b}=1$ is the $\textrm{U}(1)$ charge carried by the electron. As demonstrated in App. C, the gauge-invariant response remains independent of the specific charge assignment $(e_{f},e_{b})$. For simplicity, we can choose $(e_{f},e_{b})=(0,1)$ and refer to the bosonic parton $b$ as the chargon.
The next step is to go to the dual vortex representation of the chargon sector
$\mathcal{L}=\mathcal{L}_{\textrm{FS}}[f,a]+\frac{\mathtt{i}}{2\pi}\tilde{a}\wedge\textrm{d}(A-a)+\mathcal{L}_{\textrm{vrtx}}[v,\tilde{a}]+\ldots$ (19)
where the flux of the gauge field $\tilde{a}$ represents the density of $b$. It is important to note that under gauge constraint, the density of $b$ equals to the density of $f$, as well as the density of electrons. We are interested in the parton mean-field state (before turning on gauge-field fluctuations), where the $f$-fermions occupy the same Fermi-surface state as the original electrons. The Landau FL phase of electrons can be reproduced when the vortices $v$ are trivially gapped. In this case, the Maxwell term $\frac{1}{2e^{2}}\textrm{d}\tilde{a}\wedge\star\textrm{d}\tilde{a}$ becomes important, and the equation of motion of $\tilde{a}$ leads to a mass term $\frac{e^{2}}{8\pi^{2}}(a_{\mu})^{2}$ for the gauge field $a_{\mu}$. In the IR, the Fermi-surface state of $f$-fermions essentially corresponds to the ordinary metallic phase of gauge-invariant electrons. Conversely, various interesting electronic phases can be realized by placing the vortices $v$ in different states. For instance, incompressible Mott insulators [50, 51, 54] can be realized when the vortices are in the Higgs phase or certain topological orders. Compressible CFLs [76, 55, 56] can also be realized by putting the vortices in integer QH states. Although we do not consider translation symmetry breaking in this paper, density-wave states [77, 54, 56] can also be conveniently described by the theory in Eq. (19), where the vortex band structure has multiple minima in the Brillouin zone, and vortex condensation results in lattice translation symmetry breaking.
Central to our discussion is the electromagnetic response of the system. By definition, the conserved U(1) current in the vortex theory Eq. (19) is given by
$J^{\mu}=\frac{\delta\mathcal{L}}{\delta A_{\mu}}=\frac{\mathtt{i}}{2\pi}\varepsilon^{\mu\nu\rho}\partial_{\nu}\tilde{a}_{\rho}.$ (20)
Therefore, the response function $\Pi^{\mu\nu}=\langle J^{\mu}J^{\nu}\rangle$ is determined by the fully dressed gauge-field propagator $\langle\tilde{a}_{\mu}\tilde{a}_{\nu}\rangle$. For the problem of a Fermi surface coupled to a gauge field $a_{\mu}=(a_{\tau},\bm{a})$, it is convenient to use the Coulomb gauge $\nabla\cdot\bm{a}=0$ [30, 31, 63, 78, 34]. At the level of the random phase approximation (RPA), Eq. (19) leads to the effective theory
$\mathcal{S}=\int_{k}\tilde{a}(-k)\frac{-\tilde{\Pi}(k)}{2}\tilde{a}(k)-\tilde{a}(-k)\Pi_{\textrm{CS}}(k)A(k)+\ldots$ (21)
where $k=(\omega,\bm{k})$ is a collective notion of frequency and momentum. The response kernels $\tilde{\Pi}$ and $\Pi_{\textrm{CS}}$ are $2\times 2$ matrices in the basis $\tilde{a}=(\tilde{a}_{\tau},\tilde{a}_{T})$ and $A=(A_{\tau},A_{T})$. In our convention (see App. B), the Chern-Simons kernel is $\Pi_{\textrm{CS}}(k)=\frac{|\bm{k}|}{2\pi}\sigma^{1}$, where $\sigma^{1}$ denotes the first Pauli matrix. The fully dressed $\tilde{\Pi}(k)$ contains contributions from both the fermion sector and the vortex sector
$\tilde{\Pi}=-\Pi_{\textrm{CS}}\Pi_{f}^{-1}\Pi_{\textrm{CS}}+\Pi_{v},$ (22)
where $\Pi_{f}$ is the response of the fermionic partons $f$ to $a_{\mu}$ and $\Pi_{v}$ is the response of the vortices $v$ to $\tilde{a}_{\mu}$. The electromagnetic response of electrons is then given by
$\Pi^{-1}=-\Pi_{\textrm{CS}}^{-1}\tilde{\Pi}\Pi_{\textrm{CS}}^{-1}=\Pi_{f}^{-1}-\Pi_{\textrm{CS}}^{-1}\Pi_{v}\Pi_{\textrm{CS}}^{-1}.$ (23)
This is nothing but the Ioffe-Larkin rule Eq. (80) for the parton construction Eq. (17), where the duality relation in the chargon sector imposes $\Pi_{v}=-\Pi_{\textrm{CS}}\Pi_{b}^{-1}\Pi_{\textrm{CS}}$. To avoid redundant terminology, we refer to the equivalent expressions under duality transformations, Eq. (80), Eq. (23), and Eq. (22), as the Ioffe-Larkin rule [79].
From a different perspective, for any fermionic systems with a global U(1) symmetry, one can always introduce a gauge field $\tilde{a}_{\mu}$ to represent the conserved current $J^{\mu}$ via Eq. (20). Consequently, Eq. (21) can be interpreted as the bosonization [80, 81] of non-relativistic electrons.
Within the vortex theory framework, the charge disorder operator is represented by the spatial Wilson loop operator of the gauge field $\tilde{a}$
$\mathcal{W}_{\mathcal{C}}(\vartheta)=\exp\left(\frac{\mathtt{i}\vartheta}{2\pi}\int_{\mathcal{C}}\tilde{a}\right),$ (24)
where $\mathcal{C}$ is a closed loop in real space, and $\vartheta$ is a real-valued parameter. The expectation value $\langle\mathcal{W}_{\mathcal{C}}(\vartheta)\rangle$ can be viewed as a generating functional, with its 2nd cumulant giving the bipartite fluctuations defined in Eq. (1)
$\mathcal{F}_{\mathcal{A}}=\lim_{\vartheta\rightarrow 0}(-\mathtt{i}\partial_{\vartheta})^{2}\log\langle\mathcal{W}_{\mathcal{C}}(\vartheta)\rangle,$ (25)
where $\mathcal{C}=\partial\mathcal{A}$. The large-scale scaling of $\mathcal{F}_{\mathcal{A}}$ is determined by the gauge-field propagator $\langle\tilde{a}_{\mu}\tilde{a}_{\nu}\rangle$ through the Ioffe-Larkin rule Eq. (22).
In evaluating Eq. (25), one must be very careful and choose a gauge-invariant regularization scheme to handle the UV divergence. For example, directly using the expression in Eq. (22) (under the Coulomb gauge) and setting a real-space cut-off on the integration interval along $\mathcal{C}$ would spoil gauge invariance. We leave the discussion of these technical details to App. D.
(1) FL-CFL transition [55, 56] (2) Mott transition [50, 51] (3) Mott transition [54] $\ldots$
electron $c$ spinless $c(\bm{x})=b(\bm{x})f(\bm{x})$ spinful $c_{\sigma}(\bm{x})=b(\bm{x})f_{\sigma}(\bm{x})$ spinful $c_{\sigma}(\bm{x})=b_{\sigma}(\bm{x})f_{\sigma}(\bm{x})$ldots
fermion $f$ Fermi surface Fermi surface Fermi surface Fermi surface
boson $b$ superfluid to Laughlin at $\nu=1/2$ superfluid to trivial gapped superfluid to $Z_{N}$ topological order superfluid to …
vortex $v$ trivial gapped to IQH at $\nu=-2$ trivial gapped to Higgs trivial gapped to Higgs $\textrm{U}(1)\rightarrow Z_{N}$ trivial to …
Table 3: The parton constructions $c\sim bf$ for the quantum phase transitions out of spinless and spin-$1/2$ Landau Fermi liquids at half-filling, where $\sigma=\uparrow,\downarrow$ represents the spin quantum number. The dual vortex $v$ is introduced in Eq. (19).
IV Continuous CFL-FL Transition
We begin by discussing the critical Fermi surfaces for spinless electrons, focusing on the continuous transition between a Landau FL and a CFL at half-filling. The universal field theory of this transition was first explored in Ref. [55] and later refined in Ref. [56], with particular attention to its potential realization in twisted $\textrm{MoTe}_{2}$ bilayers through the tuning of the electron bandwidth. More recently, Ref. [71] reported that a CFL-FL transition is indeed realized in a microscopic model of spatially modulated Landau levels by adjusting the interaction strength, as evidenced by the vanishing and nonvanishing values of the Drude weight, although whether the transition is continuous remains to be confirmed.
In Sec. IV.1, we briefly discuss the vortex theory that connects Landau FL theory with HLR theory [63] and its relation to previous constructions [55, 56] through fermionic particle-vortex duality. In Sec. IV.2, we clarify the prediction of a universal constant term in Eq. (12) within the CFL phase, as derived from HLR theory, and the universal logarithmic term in Eq. (16) at the critical point, further examining its connection to the critical resistivity jump described in Eq. (15).
IV.1 Critical Theories
Within the vortex theory framework, the critical theory can be described by Eq. (19), where the fermionic vortices undergo a Chern number changing transition from $\mathsf{C}=0$ to $\mathsf{C}=-2$. This transition can be formulated using two Dirac fermions
[…]
where $\textrm{D}_{\tilde{a}}=\partial-\mathtt{i}\tilde{a}$ denotes the gauge covariant derivative. Note that each Dirac fermion is defined through the Pauli-Villars scheme with another heavy Dirac fermion in the UV. There are two phases depending on the sign of the fermion mass term $m(\bar{\psi}_{1}\psi_{1}+\bar{\psi}_{2}\psi_{2})$. In the case of $\mathsf{C}=0$, as argued in Sec. III, the Maxwell term in Eq. (26) becomes important and causes the system to flow back to the Landau FL phase of electrons. In the case of $\mathsf{C}=-2$, the vortex sector in Eq. (19) is given by $\mathcal{L}_{\textrm{vrtx}}=\frac{2\mathtt{i}}{4\pi}\tilde{a}\wedge\textrm{d}\tilde{a}$. After integrating out the gauge field $\tilde{a}$, we have the HLR theory [63] for half-filled Landau level
$\mathcal{L}=\mathcal{L}_{\textrm{FS}}[f,a]-\frac{1}{2}\frac{\mathtt{i}}{4\pi}(a-A)\wedge\textrm{d}(a-A),$ (27)
where each $f$-fermion becomes a composite fermion with two flux quanta attached.
For completeness, let us briefly mention the critical theory discussed in Ref. [55, 56], and illustrate its relation to the vortex theory Eq. (19) together with Eq. (26). The starting point is the parton construction Eq. (17) described in Sec. III. The bosonic parton $b$ is assumed to be further fractionalized into two fermions
$b(\bm{x})=f_{1}(\bm{x})f_{2}(\bm{x})$ (28)
introducing an additional $\textrm{U}(1)$ gauge field denoted by $\widehat{a}$. One can assign the charge of $A-a$ to $f_{2}$, and put it in a mean-field state with Chern number $\mathsf{C}_{2}=1$. Then a Chern number changing phase transition of $f_{1}$ from $\mathsf{C}_{1}=-1$ to $\mathsf{C}_{1}=1$ will drive the transition of the chargon $b$ from a superfluid state to the bosonic Laughlin state at $\nu=1/2$. The band touching of $f_{1}$ typically involves two massless Dirac fermions $\chi_{1},\chi_{2}$. The full critical theory can be written as
[…]
$-\frac{\mathtt{i}}{2\pi}\widehat{a}\wedge\textrm{d}(A-a)-\frac{\mathtt{i}}{4\pi}(A-a)\wedge\textrm{d}(A-a),$ (29)
where the transition is driven by the fermion mass term $m(\bar{\chi}_{1}\chi_{1}+\bar{\chi}_{2}\chi_{2})$. An important insight from Ref. [56] is that a direct second-order phase transition can be protected by translation symmetry and filling constraints.
As we show in App. E, the critical theories described by Eq. (29) and Eq. (19) (together with Eq. (26)) can be related through fermionic particle-vortex duality [82, 83], where $\psi_{1},\psi_{2}$ can be interpreted as the fermionic vortices of $\chi_{1},\chi_{2}$. Consequently, the transition from the superfluid to the Laughlin state in the chargon sector can be effectively described by the integer QH transition of vortices. The analysis presented in Ref. [56] can be carried over for Eq. (26), demonstrating that other fermion bilinears are not permitted by translation symmetry under filling constraints.
IV.2 Charge Fluctuations
Our primary focus is on understanding the static structure factor Eq. (8) at long wavelengths, as it governs the leading-order scaling of bipartite fluctuations $\mathcal{F}$ (see Tab. 1 and App. D). The subleading corner contribution in Eq. (12) or Eq. (16) depends on whether the localization tensor $\mathsf{G}^{ij}$, as described in Eq. (9) and Eq. (10), is finite or logarithmically divergent. In Landau FLs, it is well understood that the leading-order scaling follows $\mathcal{F}\sim L\log(L)$, while the subleading corner term is not well-defined due to the slow decay of spatial charge correlations [21]. This non-local feature of Landau FLs is tied to the power-law divergence of $\mathsf{G}^{ij}$. In the following, we will begin our analysis with the CFL phase and then proceed to examine the critical point.
IV.2.1 CFL Phase
In this subsection, we consider the standard RPA analysis of the HLR theory Eq. (27). For simplicity, we primarily focus on cases with continuous translational and rotational symmetries, as in Refs. [63, 34], with comments on the anisotropic cases provided later. Under the Coulomb gauge, the response function of the $f$-fermions is expressed as $\Pi_{f}(k)=\textrm{diag}(\Pi_{f}^{\tau\tau}(k),\Pi_{f}^{TT}(k))$, where $k=(\omega,\bm{k})$. At the one-loop level, the longitudinal and transverse components are
$\Pi_{f}^{\tau\tau}(k)$ $=-\mathscr{D}_{F}\left(1-\frac{|\omega|}{\sqrt{\omega^{2}+(v_{F}\bm{k})^{2}}}\right),$
$\Pi_{f}^{TT}(k)$ $=\mathscr{D}_{F}\frac{|\omega|\sqrt{\omega^{2}+(v_{F}\bm{k})^{2}}-\omega^{2}}{|\bm{k}|^{2}}.$ (30)
Here, $v_{F}$ is the fermi velocity, $k_{F}$ denotes the fermi momentum, and $\mathscr{D}_{F}=\frac{k_{F}}{2\pi v_{F}}$ represents the density of states at the fermi level. The vortex sector is in the integer QH state at the filling $\nu=-2$, which has the response function $\Pi_{v}(k)=-2\Pi_{\textrm{CS}}(k)=-2\frac{|\bm{k}|}{2\pi}\sigma^{1}$. The electron density-density correlation is determined using the Ioffe-Larkin rule Eq. (23). Upon integrating over all frequencies $\omega$, the static structure factor is obtained as follows
$\Pi^{\tau\tau}(\tau\rightarrow 0,\bm{k})=\int\frac{\textrm{d}\omega}{2\pi}\frac{1}{\Pi_{f}^{\tau\tau}(k)^{-1}-\frac{(4\pi)^{2}}{|\bm{k}|^{2}}\Pi_{f}^{TT}(k)}$
$\approx\frac{1-e^{-(|\bm{k}|\ell_{B})^{2}/2}}{4\pi\ell_{B}^{2}}-\frac{|\bm{k}|^{3}\log(1/|\bm{k}|)}{4\pi^{2}k_{F}}+\ldots$ (31)
where $\ell_{B}$ denotes the magnetic length scale.
The first term in Eq. (31) arises from the inter-Landau-level gapped modes [84, 85]. Based on calculations similar to those in Ref. [21], we find a corner contribution of $-\textrm{f}(\theta)/(8\pi^{2})$ to the bipartite charge fluctuations, where $\textrm{f}(\theta)$ is the universal angle function defined in Eq. (6). The second term $|\bm{k}|^{3}\log|\bm{k}|$ arises from the gapless modes in the lowest Landau level. This long-wavelength result has been predicted by both the HLR theory [63, 86] as well as the Son-Dirac theory [82, 87]. As we illustrate in App. D, the gapless modes contribute only to a boundary-law term. By combining the contributions from both gapped and gapless modes, we obtain the final result in Eq. (7) for the configuration depicted in Fig. 1. It is noteworthy that Eq. (7) applies to both the mentioned examples of gapped QH insulators [21] and gapless CFLs.
It would be highly compelling if Eq. (7) holds true for all charge insulators, i.e., if the corner contribution is universally determined by the DC Hall conductivity $\sigma_{xy}$. However, unlike at the critical point, deep inside the insulator phase, the gapped modes exhibit short-range correlations and may be sensitive to microscopic details. As we clarify in a follow-up work [75], for systems with continuous translational and rotational symmetries, Eq. (12) generally holds, but with a non-universal value of the localization tensor $\mathsf{G}\geq|\sigma_{xy}|/2$. In other words, the coefficient of the universal angle function satisfies a universal lower bound set by the topological invariant $\sigma_{xy}$. In Ref. [75], we explain this bound and its saturation from the perspective of many-body quantum geometry [^4] Here, $\mathsf{G}^{ij}=\mathsf{G}\delta^{ij}$ serves as the many-body quantum metric [75], defined by adiabatic flux insertion (or twisted boundary conditions), where $i=x,y$ denotes the spatial indices.. It turns out that $\mathsf{G}=|\sigma_{xy}|/2$ is guaranteed under specific conditions, such as Galilean invariance and wavefunction holomorphicity, which are indeed satisfied by many examples. In fact, our predicted formula Eq. (7) has already been numerically verified by Monte Carlo simulations in Ref. [88] using the Rezayi-Read wavefunction, considering CFLs for fermions at $\nu=1/2$ and $\nu=1/4$, as well as CFLs for bosons at $\nu=1$ and $\nu=1/3$ fillings.
More generally, the system may lack continuous translational and rotational symmetries, as in the models with spatially modulated periodic potentials discussed in Ref. [57, 58, 71], which concern the half-filled Chern band in Morié materials. In such cases, the corner contribution may depend on the corner angle, subregion orientation, and other microscopic details. Nevertheless, in Ref. [75], we demonstrate that the corner contribution can be expressed as an infinite series of orientation-resolved universal angle functions, generalizing the formula Eq. (6), along with their non-universal coefficients. However, these details are beyond the scope of this work.
IV.2.2 Critical Point
Next, we turn our attention to the critical point, where the analysis becomes slightly more intricate. The QH transition in the vortex sector (i.e., the chargon sector) is presumably described by a CFT. It was argued in Ref. [55, 56] that the CFT sector and the $f$-fermion sector are dynamically decoupled right at the critical point. This assertion is supported by two key observations. (1) The Landau damping term of the gauge field $a_{\mu}$ effectively suppresses gauge fluctuations at the critical point. (2) When considering an operator $O$ from the CFT sector coupled to the $f$-fermion via $Of^{\dagger}f$, an additional Landau damping term $\int_{k}\frac{|\omega|}{|\bm{k}|}|O(k)|$ is generated. However, this term becomes irrelevant if the scaling dimension satisfies $\Delta[O]>3/2$, a condition likely met in the critical theory under consideration [55, 56].
Assuming the dynamical decoupling of the two sectors, the critical response in the vortex/chargon sector still maintains the CFT form [^5] In terms of space-time components, the response function has the form $\Pi_{\textrm{CFT}}^{\mu\nu}(k)=-\sigma_{xx}|k|(\delta_{\mu\nu}-\frac{k_{\mu}k_{\nu}}{|k|^{2}})-\sigma_{xy}\varepsilon_{\mu\nu\rho}k_{\rho}$.. Under the Coulomb gauge, it is expressed as
$\Pi_{\textrm{CFT}}(k)=\left(\begin{array}[]{cc}-\frac{|\bm{k}|^{2}}{\omega}\mathsf{F}_{L}(\frac{\omega}{|\bm{k}|})&-\sigma_{xy}|\bm{k}|\\ -\sigma_{xy}|\bm{k}|&\omega\mathsf{F}_{T}(\frac{\omega}{|\bm{k}|})\end{array}\right),$ (34)
where the dimensionless scaling functions of the longitudinal and transverse components are given by
$\mathsf{F}_{L}(\lambda)$ $=\sigma_{xx}\frac{\lambda}{\sqrt{1+\lambda^{2}}},$
$\mathsf{F}_{T}(\lambda)$ $=\sigma_{xx}\frac{\sqrt{1+\lambda^{2}}}{\lambda}.$ (35)
Here, $\sigma_{xx}$ and $\sigma_{xy}$ represent the universal longitudinal and Hall conductivities of the CFT. Their values in $\Pi_{v}(k)$ for vortices and in $\Pi_{b}(k)$ for chargons are related by
$\sigma_{xx}^{b}$ $=\frac{1}{(2\pi)^{2}}\frac{\sigma_{xx}^{v}}{(\sigma_{xx}^{v})^{2}+(\sigma_{xy}^{v})^{2}},$
$\sigma_{xy}^{b}$ $=\frac{1}{(2\pi)^{2}}\frac{-\sigma_{xy}^{v}}{(\sigma_{xx}^{v})^{2}+(\sigma_{xy}^{v})^{2}}.$ (36)
According to the Ioffe-Larkin composition rule Eq. (23) (or Eq. (80)), the static structure factor of electrons can be calculated as follows
$\Pi^{\tau\tau}(\tau\rightarrow 0,\bm{k})=\int\frac{\textrm{d}\omega}{2\pi}\frac{|\bm{k}|^{2}\Pi_{f}^{\tau\tau}(k)(\mathsf{F}_{L}^{b}(\frac{\omega}{|\bm{k}|})(\omega\mathsf{F}_{T}^{b}(\frac{\omega}{|\bm{k}|})+\Pi_{f}^{TT}(k))-\omega(\sigma_{xy}^{b})^{2})}{(\omega\mathsf{F}_{T}^{b}(\frac{\omega}{|\bm{k}|})+\Pi_{f}^{TT}(k))(|\bm{k}|^{2}\mathsf{F}_{L}^{b}(\frac{\omega}{|\bm{k}|})-\omega\Pi_{f}^{\tau\tau}(k))-\omega|\bm{k}|^{2}(\sigma_{xy}^{b})^{2}}=-\frac{\sigma_{xx}^{b}}{\pi}|\bm{k}|^{2}\log(1/|\bm{k}|)+\ldots$ (37)
Here, $\Pi_{f}^{\tau\tau}(k)$ and $\Pi_{f}^{TT}(k)$ are the components of the response function of $f$-fermions given in Eq. (30). The scaling functions $\mathsf{F}_{L}^{b}(k)$ and $\mathsf{F}_{T}^{b}(k)$ are provided by Eq. (35), with the coefficient $\sigma_{xx}^{b}$. Evaluating the frequency integral in Eq. (37) is not an easy task. In App. F, we present two methods for extracting the long-wavelength behavior $|\bm{k}|^{2}\log(1/|\bm{k}|)$. By retaining only the leading-order term of the integrand in the small-$\bm{k}$ expansion, the $\omega$-integral can be performed analytically. Additionally, we have numerically evaluated the full expression in Eq. (37). Remarkably, at small $|\bm{k}|/k_{F}$, the numerical result aligns closely with the analytical expression.
The Fourier transform of Eq. (37) yields the power-law spatial correlation given by Eq. (2), with an exponent $\alpha=4$. The overall coefficient $C_{0}$ is given by the universal constant $C_{\rho}=2(\sigma_{xx}^{b})/\pi^{2}$, which identifies the current central charge of the CFT describing the chargon sector. By examining the integrand in Eq. (37) and sequentially taking the limits $\omega\rightarrow 0$ and then $\bm{k}\rightarrow 0$, we can demonstrate the vanishing compressibility. Moreover, the vanishing Drude weight can be confirmed using the Ioffe-Larkin rule Eq. (23) (or Eq. (80)). As a result, we identify an incompressible state with vanishing Drude weight, characterized by a logarithmically divergent localization tensor as described by Eq. (14), despite the absence of conformal symmetry.
According to Refs. [20, 18, 19, 21] (also see App. D), for the subregion depicted in Fig. 1, the bipartite charge fluctuations follow Eq. (16). Similar to the first term in Eq. (31), note that Eq. (37) also includes a term $|\bm{k}|^{2}$ arising from gapped modes, meaning that the localization tensor contains subleading contributions in addition to the divergence described in Eq. (14). These modes contribute to a corner contribution $(\textrm{const})\textrm{f}(\theta)$ that does not scale with $L$, similar to the subleading term in Eq. (12). The coefficient of this term is a function of $\sigma_{xx}^{b}$ and $\sigma_{xy}^{b}$. However, isolating this subleading constant term from the $\log(L)$ scaling in experiments or numerical simulations can be challenging. Additionally, in theory, this constant term is also sensitive to the UV cutoff.
One may wonder whether the universal number $C_{\rho}$ can be detected through other experimental observables, such as transport measurements. A promising candidate is the predicted resistivity jump at the critical point [56], which is solely determined by the universal data of the CFT sector. The key argument leading to the universal resistivity jump is based on the observation that $\Pi^{\tau j}(\mathtt{i}\omega,\bm{k}\rightarrow 0)=0$, where $j=x,y$, holds for both the $f$-fermion sector and the $b$-chargon sector. Considering the spatial components of the Ioffe-Larkin rule Eq. (23) (or Eq. (80)), the total resistivity is given by $\rho_{ij}=\rho^{f}_{ij}+\rho^{b}_{ij}$. As one approaches the critical point from the Landau FL phase, additional contributions to the longitudinal and Hall components arise from the gapless degrees of freedom in the CFT:
$\Delta\rho_{xx}$ $=\rho_{xx}^{b}=\frac{\sigma_{xx}^{b}}{(\sigma_{xx}^{b})^{2}+(\sigma_{xy}^{b})^{2}},$
$\Delta\rho_{xy}$ $=\rho_{xy}^{b}=\frac{-\sigma_{xy}^{b}}{(\sigma_{xx}^{b})^{2}+(\sigma_{xy}^{b})^{2}}.$ (38)
These equations allow us to solve for $\sigma_{xx}^{b}$ and express the universal coefficient $C_{\rho}=2(\sigma_{xx}^{b})/\pi^{2}$ in Eq. (16) as a function of $\Delta\rho_{xx}$ and $\Delta\rho_{xy}$, as in Eq. (15). Note that both $\Delta\rho_{xx}$ and $\Delta\rho_{xy}$ are scaling functions of $\omega/T$ (i.e., frequency over temperature), and Eq. (15) should be understood as a relation in the limit $T\rightarrow 0$. The values of $\Delta\rho_{xx}$ and $\Delta\rho_{xy}$, obtained through the large-$N$ expansion, are available in Refs. [89, 76], and could potentially be determined more precisely using the numerical method of conformal bootstrap [64].
V Continuous Mott Transition
[…]
V.1 Charge Fluctuations
To cause the Fermi surface of electrons to disappear abruptly in a continuous manner, spin-charge separation is necessary, leaving a neutral Fermi surface on the insulator side. The original theoretical proposal [49, 50, 51] was based on the parton construction
$c_{\sigma}(\bm{x})=b(\bm{x})f_{\sigma}(\bm{x}),$ (39)
where each electron $c_{\sigma}$ is fractionalized into a spinless bosonic chargon $b$, which carries the electric charge, and a charge-neutral fermionic spinon $f_{\sigma}$, which carries the spin quantum number. There is a dynamical $\textrm{U}(1)$ gauge field $a_{\mu}=(a_{\tau},\bm{a})$ that couples $b$ and $f$. After introducing the dual bosonic vortices $\varphi$ of the chargons $b$, the critical theory is described by Eq. (19), where $\mathcal{L}_{\textrm{FS}}[f_{\sigma},a]$ represents the Fermi-surface state of spinons. The continuous Mott transition is triggered by the superfluid to Mott insulator transition of chargons. In the dual picture, the vortex condensation is described by
$\mathcal{L}_{\textrm{vrtx}}[\varphi,\tilde{a}]=|\textrm{D}_{\tilde{a}}\varphi|^{2}+r|\varphi|^{2}+u|\varphi|^{4},$ (40)
where $\textrm{D}_{\tilde{a}}=\partial-\mathtt{i}\tilde{a}$ denotes the gauge covariant derivative. Here, a single parameter $r\sim(g-g_{c})$ controls the transition, with $g$ denoting the bandwidth of the original electrons. When $r>0$, the vortices $\varphi$ are trivially gapped, resulting in the Landau FL phase of electrons. Conversely, when $r<0$, the condensation of vortices Higgs the gauge field $\tilde{a}$, leading the system to become a spin-liquid Mott insulator with gapless spinons $f_{\sigma}$.
An important feature of the critical theory is that the chargon/vortex sector is dynamically decoupled from the spinon sector, as the conditions outlined at the beginning of Sec. IV.2.2 are indeed satisfied by the XY fixed point [50, 51]. As a result, the zero-temperature response function $\Pi_{b}$ of the chargons follows the CFT form given by Eq. (34), where $\sigma_{xx}^{b}$ represents the critical conductivity of the 3D XY universality class, and the Hall conductivity $\sigma_{xy}^{b}$ vanishes. In the dual picture, the response function of vortices $\Pi_{v}$ is also described by Eq. (34) with the coefficients $\sigma_{xx}^{v}=1/(4\pi^{2}\sigma_{xx}^{b})$ and $\sigma_{xy}^{v}=0$. Assuming the $f$-fermions remain in the mean-field state with the response function $\Pi_{f}$ as given by Eq. (30), we can apply the Ioffe-Larkin rule Eq. (23) (or Eq. (80)) to derive the leading-order long-wavelength behavior of the static structure factor
$\Pi^{\tau\tau}(\tau\rightarrow 0,\bm{k})$ $=\int\frac{\textrm{d}\omega}{2\pi}\frac{1}{\Pi_{f}^{\tau\tau}(k)^{-1}-\frac{\omega}{|\bm{k}|^{2}}\mathsf{F}^{b}_{L}(\frac{\omega}{|\bm{k}|})^{-1}}$
$=-\frac{\sigma_{xx}^{b}}{\pi}|\bm{k}|^{2}\log(1/|\bm{k}|)+\ldots$ (41)
Since the Hall response vanishes, Eq. (41) for the Mott transition is much simpler than Eq. (37) for the CFL-FL transition. The common feature of these two types of critical points is that they correspond to incompressible states with vanishing Drude weight, where the static structure factor exhibits a CFT-like behavior, with the overall coefficient determined by the universal conductivity $\sigma_{xx}^{b}$ of the chargons.
In this case, the bipartite charge fluctuations are described by Eq. (16) for the configuration in Fig. 1, with a universal corner contribution proportional to
$C_{\rho}=C_{J}^{(\textrm{XY})}$ (42)
where $C_{J}^{(\textrm{XY})}=2\sigma_{xx}^{b}/\pi^{2}$ represents the current central charge of the 3D XY fixed point. A notable property arising from time-reversal invariance is the direct relationship between the universal corner contribution and the longitudinal resistivity jump at the critical point. With a vanishing Hall response, the total longitudinal resistivity is given by the Ioffe-Larkin rule $\rho_{xx}=\rho_{xx}^{f}+\rho_{xx}^{b}$, where $\rho_{xx}^{f}$ is the spinon contribution and $\rho_{xx}^{b}$ is the chargon contribution. At $T=0$, in the insulating phase, the gapped chargon sector leads to $\rho_{xx}^{b}=+\infty$, resulting in infinite total electrical resistivity $\rho_{xx}$. In the metallic phase, the chargon superfluid has $\rho_{xx}^{b}=0$, so the total resistivity $\rho_{xx}$ is determined solely by the spinon contribution. At the critical point, a universal jump $\Delta\rho_{xx}(\omega/T)$ is predicted [51] due to the universal chargon contribution. Under the limit $\omega/T\rightarrow\infty$, $\Delta\rho_{xx}=2/(\pi^{2}C_{J}^{(\textrm{XY})})$ is directly related to the universal corner contribution to bipartite charge fluctuations. In the absence of disorder, such that the spinon contribution $\rho_{xx}^{f}\approx 0$, the total critical resistivity can be approximated as $\Delta\rho_{xx}$.
V.2 Spin Fluctuations
In addition to the $\textrm{U}(1)$ charge symmetry, the system also possesses a $\textrm{U}(1)$ spin symmetry associated with the conservation of the third component of spin. A deeper understanding can be achieved by incorporating both charge and spin vortices $\varphi^{c},\varphi^{s}$ in strongly correlated metals. The full theory is introduced as follows
$\mathcal{L}=\;$ $\mathcal{L}_{\textrm{FS}}[f,a^{c},a^{s}]+\frac{\mathtt{i}}{2\pi}\tilde{a}^{c}\wedge\textrm{d}(A^{c}-a^{c})+\frac{\mathtt{i}}{2\pi}\tilde{a}^{s}\wedge\textrm{d}(A^{s}-a^{s})+\frac{1}{2e_{c}^{2}}\textrm{d}\tilde{a}^{c}\wedge\star\textrm{d}\tilde{a}^{c}+\frac{1}{2e_{s}^{2}}\textrm{d}\tilde{a}^{s}\wedge\star\textrm{d}\tilde{a}^{s}$
$+|(\partial-\mathtt{i}\tilde{a}^{c})\varphi^{c}|^{2}+r_{c}|\varphi^{c}|^{2}+u_{c}|\varphi^{c}|^{4}+|(\partial-\mathtt{i}\tilde{a}^{s})\varphi^{s}|^{2}+r_{s}|\varphi^{s}|^{2}+u_{s}|\varphi^{s}|^{4}+\ldots.$ (43)
Here, the background fields $A^{c}$ and $A^{s}$ are introduced to keep track of the two $\textrm{U}(1)$ symmetries. The spinon $f=(f_{\uparrow},f_{\downarrow})$ still forms a Fermi-surface state and is coupled to emergent $\textrm{U}(1)$ gauge fields $a^{c}$ and $a^{s}$ through the covariant derivative $\textrm{D}=\partial\sigma^{0}-\mathtt{i}a^{c}\sigma^{0}-\mathtt{i}a^{s}\sigma^{3}$, where $\sigma^{\mu}$ denotes the vector of Pauli matrices. The fluxes of the $\textrm{U}(1)$ gauge fields $\tilde{a}^{c}$ and $\tilde{a}^{s}$ represent the densities of the bosonic partons that carry the charge-$\textrm{U}(1)$ and spin-$\textrm{U}(1)$ symmetries. The parton construction for Eq. (43) resembles that of the $\textrm{U}(2)$ gauge theory discussed in Ref. [91]. There is no prior relationship between the two sets of coupling constants $e,r,u$ for the different vortices.
Based on the vortex theory Eq. (43), two gauge-invariant Wilson loop operators can be defined:
$\mathcal{W}_{\mathcal{C}}^{c}(\vartheta)$ $=\exp\left(\frac{\mathtt{i}\vartheta}{2\pi}\int_{\mathcal{C}}\tilde{a}^{c}\right),$
$\mathcal{W}_{\mathcal{C}}^{s}(\vartheta)$ $=\exp\left(\frac{\mathtt{i}\vartheta}{2\pi}\int_{\mathcal{C}}\tilde{a}^{s}\right),$ (44)
which represent the charge-$\textrm{U}(1)$ and spin-$\textrm{U}(1)$ disorder operators associated with the subsystem $\mathcal{A}$, where $\mathcal{C}=\partial\mathcal{A}$. The charge and spin fluctuations $\mathcal{F}_{\mathcal{A}}^{c}$ and $\mathcal{F}_{\mathcal{A}}^{s}$ are then obtained using $\mathcal{W}_{\mathcal{C}}^{c}$ and $\mathcal{W}_{\mathcal{C}}^{s}$, as described in Eq. (25).
The vortex theory described by Eq. (43) is a multi-critical theory governed by two coupling constants, $r_{c}$ and $r_{s}$. The Mott transition discussed earlier corresponds to the scenario where the spin vortex $\varphi^{s}$ is gapped, while the condensation of the charge vortex $\varphi^{c}$ is dual to the XY transition. When both $\varphi^{c}$ and $\varphi^{s}$ are gapped, the system remains in the ordinary metallic phase. However, if $\varphi^{c}$ condenses, we are left with gapless spinons coupled to a deconfined gauge field $a^{c}$, while the gauge field $\tilde{a}^{c}$ is gapped out due to the Higgs mechanism. This is the Mott insulator with a spinon Fermi surface. Conversely, if $\varphi^{c}$ is gapped and $\varphi^{s}$ condenses, it leads to a transition from an ordinary metal to an algebraic charge liquid with power-law charge correlations. The behaviors of the charge and spin bipartite fluctuations are summarized in Tab. 4. Regarding the final fate in the IR, there is an important difference between the vortex-condensed phases of $\varphi^{c}$ and $\varphi^{s}$. In the case of $\varphi^{s}$, the gauge field $a^{s}$ will eventually induce a pairing instability in the Fermi-surface state, resulting in a ground state of a charge superconductor. Nevertheless, the charge liquid should still be observable within a finite energy window.
$\varphi^{s}$-gapped $\varphi^{s}$-condensed
$\varphi^{c}$-gapped $(L\log L,L\log L)$ $(L\log L,L)$
$\varphi^{c}$-condensed $(L,L\log L)$ $(L,L)$
Table 4: Leading-order scalings of charge and spin bipartite fluctuations $(\mathcal{F}^{c},\mathcal{F}^{s})$ across the transitions described by the multicritical vortex theory Eq. (43) involving charge and spin vortices $\varphi^{c}$ and $\varphi^{s}$. At the critical point for each vortex condensation, the corresponding bipartite fluctuations exhibit a universal corner contribution proportional to the critical conductivity associated with the 3D XY universality class.
When both $\varphi^{c}$ and $\varphi^{s}$ are condensed, the $f$-fermions become coupled to two $\textrm{U}(1)$ gauge fields
$\mathcal{L}=\mathcal{L}_{\textrm{FS}}[f,a^{c},a^{s}]+\frac{1}{2e_{c}^{2}}\textrm{d}a^{c}\wedge\star\textrm{d}a^{c}+\frac{1}{2e_{s}^{2}}\textrm{d}a^{s}\wedge\star\textrm{d}a^{s}.$ (45)
There is a competition between the effects of $a^{s}$, which drives a pairing instability, and $a^{c}$, which suppresses the tendency towards pairing. A perturbative RG analysis in Ref. [91] shows that the IR fate of the system depends on the UV coupling constants. If $e_{c}/e_{s}<1$, the Fermi-surface state becomes unstable and favors pairing. Conversely, if $e_{c}/e_{s}>1$, the system flows to a stable non-Fermi liquid fixed point. The case where $e_{c}/e_{s}=1$ resembles the behavior of Landau FLs, with the stability of the Fermi-surface state depending on whether the pairing interaction is repulsive or attractive in the UV.
V.3 Charge Fractionalization
Another type of vortex theory was introduced in Ref. [54], inspired by the experimental observation of a continuous Mott transition in the TMD heterobilayer $\textrm{MoTe}_{2}/\textrm{WSe}_{2}$ [52]. In this theory, two bosonic chargons $b_{\uparrow}$ and $b_{\downarrow}$ are introduced for the two spin/valley quantum numbers, defined as
$c_{\uparrow}(\bm{x})=b_{\uparrow}(\bm{x})f_{\uparrow}(\bm{x}),\quad c_{\downarrow}(\bm{x})=b_{\downarrow}(\bm{x})f_{\downarrow}(\bm{x}),$ (46)
where $f_{\sigma}$ represents the fermionic spinon. Because of time-reversal symmetry, the condensation of both $b_{\uparrow}$ and $b_{\downarrow}$ occurs simultaneously, leading to the metal-insulator transition. The novelty of this theory lies in the filling factor of the chargons. When the electron $c$ is at half-filling, both $b_{\uparrow}$ and $b_{\downarrow}$ are also at half-filling, in contrast to the integer filling of $b$ in the construction given by Eq. (39). In this case, the Lieb-Shultz-Matthis (LSM) theorem [92, 93] dictates that the Mott insulator phase of each chargon cannot be a trivial insulator. Instead, it must either be a topological order or form a density wave that spontaneously breaks the translation symmetry. In both cases, the critical point exhibits charge fractionalization and leads to anomalously large critical resistivity [54]. For simplicity, in this section, we focus on the case of topological order for illustration purposes.
We adopt the critical theory from Ref. [54] as follows
$\mathcal{L}=\;$ $\mathcal{L}_{\textrm{FS}}[f_{\uparrow},a^{\uparrow}]+\mathcal{L}_{\textrm{FS}}[f_{\downarrow},a^{\downarrow}]+\frac{\mathtt{i}}{2\pi}\tilde{a}^{\uparrow}\wedge\textrm{d}(A^{c}+A^{s}-a^{\uparrow})+\frac{\mathtt{i}}{2\pi}\tilde{a}^{\downarrow}\wedge\textrm{d}(A^{c}-A^{s}-a^{\downarrow})+\frac{1}{2e^{2}}\textrm{d}\tilde{a}^{\uparrow}\wedge\star\textrm{d}\tilde{a}^{\uparrow}$
$+\frac{1}{2e^{2}}\textrm{d}\tilde{a}^{\downarrow}\wedge\star\textrm{d}\tilde{a}^{\downarrow}+|(\partial-\mathtt{i}N\tilde{a}^{\uparrow})\varphi^{\uparrow}|^{2}+|(\partial-\mathtt{i}N\tilde{a}^{\downarrow})\varphi^{\downarrow}|^{2}+r(|\varphi^{\uparrow}|^{2}+|\varphi^{\downarrow}|^{2})+u|\varphi^{\uparrow}|^{4}+u|\varphi^{\downarrow}|^{4}+\ldots.$ (47)
The background fields $A^{c}$ and $A^{s}$ are defined in the same way as in Eq. (43). The two emergent $\textrm{U}(1)$ gauge fields $a^{\uparrow}$ and $a^{\downarrow}$ arise from the gauge redundancies in the parton construction Eq. (46). The fluxes of the $\textrm{U}(1)$ gauge fields $\tilde{a}^{\uparrow}$ and $\tilde{a}^{\downarrow}$ represent the densities of the chargons $b_{\uparrow}$ and $b_{\downarrow}$ respectively, while the vortex excitations $\varphi^{\uparrow},\varphi^{\downarrow}$ are charged under $\tilde{a}^{\uparrow}$ and $\tilde{a}^{\downarrow}$. In Eq. (47), the theory essentially contains two decoupled sectors labeled by the quantum numbers $\uparrow$ and $\downarrow$, as inter-valley couplings involve high-energy processes requiring large momentum transfer. Due to time-reversal symmetry, the two sectors are identical, meaning that the fermionic partons $f_{\uparrow},f_{\downarrow}$ are in the same mean-field state, with identical coupling constants in both sectors. Because of the $\pi$ flux background of $\tilde{a}^{\uparrow}$ (and $\tilde{a}^{\downarrow}$) per unit cell, the vortex dynamics are frustrated. Without breaking lattice translation symmetry, only $N$-vortex bound states can condense, where $N$ must be an even integer. In Eq. (47), the bosonic field $\varphi^{\uparrow}$ and $\varphi^{\downarrow}$ represent these $N$-vortex bound states.
The condensation of both $\varphi^{\uparrow}$ and $\varphi^{\downarrow}$ is controlled by a single coupling constant $r$, which is related to the bandwidth of the electrons. In the Mott insulator, each spin/valley sector forms a $\mathbb{Z}_{N}$ topological order. By following a similar analysis to that at the beginning of Sec. IV.2.2, it can be shown that the Landau damping effects are irrelevant at the 3D XY* transition. This implies that, once again, the vortex sector of each spin/valley is dynamically decoupled from the spinon Fermi surface. In both the Mott insulator and the critical point, the charge carriers are the anyons of the $\mathbb{Z}_{N}$ topological order. Each one carries the fractional electric charge $e_{*}=e/N$.
The charge and spin disorder operators are represented by the Wilson loop operators
$\mathcal{W}_{\mathcal{C}}^{c}(\vartheta)$ $=\exp\left(\frac{\mathtt{i}\vartheta}{2\pi}\int_{\mathcal{C}}(\tilde{a}^{\uparrow}+\tilde{a}^{\downarrow})\right),$
$\mathcal{W}_{\mathcal{C}}^{s}(\vartheta)$ $=\exp\left(\frac{\mathtt{i}\vartheta}{2\pi}\int_{\mathcal{C}}(\tilde{a}^{\uparrow}-\tilde{a}^{\downarrow})\right),$ (48)
which define the charge and spin bipartite fluctuations $\mathcal{F}_{\mathcal{A}}^{c}$ and $\mathcal{F}_{\mathcal{A}}^{s}$ through Eq. (25), where $\mathcal{C}=\partial\mathcal{A}$. The scaling behaviors of $\mathcal{F}^{c}$ can be easily seen to follow $L\log L$ and $L$ by using the gauge-field propagators $\langle\tilde{a}^{\uparrow}\tilde{a}^{\uparrow}\rangle$ and $\langle\tilde{a}^{\downarrow}\tilde{a}^{\downarrow}\rangle$ before and after the vortex condensation. At the critical point, an interesting universal corner contribution appears, with a coefficient that deviates from Eq. (42). The self-energy of $\tilde{a}^{\uparrow}$ is dominated by the loop corrections from $\varphi^{\uparrow}$, which is proportional to $N^{2}$. The two-point function of the conserved current $J^{\uparrow}=\frac{\mathtt{i}}{2\pi}\star\textrm{d}\tilde{a}^{\uparrow}$ at the XY* fixed point can be shown to be proportional to $1/N^{2}$ and follows the form
$\left\langle J_{\mu}^{\uparrow}(x)J_{\nu}^{\uparrow}(0)\right\rangle=\frac{C_{J}^{(\textrm{XY})}}{N^{2}}\frac{1}{|x|^{4}}\left(\delta_{\mu\nu}-\frac{2x_{\mu}x_{\nu}}{|x|^{2}}\right),$ (49)
where we still use $C_{J}^{(\textrm{XY})}$ to denote the current central charge at the XY fixed point. Identical expressions hold for $\tilde{a}^{\downarrow}$ and $J^{\downarrow}$. As a result, the charge fluctuations $\mathcal{F}^{c}$ now follow Eq. (16) with a universal coefficient
$C_{\rho}=\frac{2C_{J}^{(\textrm{XY})}}{N^{2}}.$ (50)
This value of $C_{\rho}$ can be used to determine the universal resistivity jump $\Delta\rho=N^{2}/(\pi^{2}C_{J}^{(\textrm{XY})})$, as delineated in Eq. (15). In Ref. [52], it is argued that disorder effects are weak. Neglecting the contributions to the total resistivity from the spinon Fermi surface, we approximate the total resistivity by the jump $\Delta\rho$. The factor of $N^{2}$ enhancement in $\Delta\rho$ may potentially explain the large critical resistivity observed in Ref. [52].
Another key difference compared to the Mott transition described by Eq. (43) is the distinct scaling of the bipartite spin fluctuations $\mathcal{F}^{s}$ in the Mott insulator. Because the vortex condensation Higgses both $\tilde{a}^{\uparrow}$ and $\tilde{a}^{\downarrow}$, the scaling follows a boundary law $\mathcal{F}^{s}\sim L$.
VI Discussion and Outlook
In this paper, we investigate the bipartite fluctuations of conserved quantities across interaction-driven “metal-insulator transitions” from a Landau FL to a “charge insulator” defined by localization criteria (see Sec. II.2). These charge insulators, which exhibit a vanishing Drude weight, include incompressible examples like spin-liquid Mott insulators and compressible examples like CFLs. If the transitions are continuous, we predict a logarithmic divergence of the localization tensor as the correlation length approaches infinity, as described in Eq. (14). This divergence is associated with a universal subleading corner contribution to bipartite fluctuations (see Eq. (16)), with the universal coefficient determined by the critical resistivity jump, as outlined in Eq. (15). Therefore, we establish a fluctuation-dissipation theorem-type relationship linking ground-state charge (or spin) fluctuations to transport properties. Our results, which stem from the general consequences of the Ioffe-Larkin composition rule, are based on two dynamically decoupled parton sectors at criticality: one forms a stable Fermi-surface state, while the other behaves like a CFT at the transition. These results are expected to hold at other similar phase transitions as well.
In Sec. IV, we also examine the CFL phase, where the localization tensor $\mathsf{G}^{ij}=\mathsf{G}\delta^{ij}$ is finite, and a subleading corner contribution is identified in Eq. (12). Based on the standard RPA analysis of the HLR theory, we find that the value of $\mathsf{G}$ is universal and is given by the Hall conductivity $\sigma_{xy}$ (see Eq. (13)). However, as we will clarify in an upcoming work [75], for a general (anisotropic) charge insulator, the value of $\mathsf{G}^{ij}$ is non-universal and is subject to both a topological lower bound and an energetic upper bound [60, 61, 62, 72]
00 0 $\leq\mathsf{G}^{ij}-\frac{\mathtt{i}}{2}\sigma_{xy}\epsilon^{ij}$
00 0 $\leq\lim_{\bm{A}\rightarrow 0}\frac{1}{2\Delta}\frac{1}{V}\left\langle\!\frac{\partial^{2}H}{\partial A_{i}\partial A_{j}}\!\right\rangle-\mathsf{G}^{ij}$ (51)
where the right-hand side of each line is a positive semi-definite matrix. Here, $\Delta$ denotes the energy gap of the many-body Hamiltonian $H$, $V$ is the volume of the system, and $\bm{A}$ represents the background U(1) field. The specific conditions for the saturation of these bounds will be further elaborated in Ref. [75].
In Sec. IV and Sec. V, we perform our calculations of static structure factors at the critical points (e.g., Eq. (37) and Eq. (41)) within the RPA. We expect these results to hold as long as fluctuation effects are not too strong. The Ioffe-Larkin rule $\Pi^{-1}=\Pi_{f}^{-1}+\Pi_{b}^{-1}$, arising from local gauge constraints, is expected to hold at the level of linear response [78]. Regarding the expression of $\Pi_{f}$, as discussed in Ref. [30, 31, 34, 33, 36, 41], the charge susceptibility of the Fermi surface does not contain any singular contributions. We anticipate that the long-wavelength behavior of $\Pi^{\tau\tau}$ is still dominated by $\Pi_{b}$ in the CFT sector, which has a rigid structure described by Eq. (34). To support this conclusion, one could attempt to derive the full expression of $\Pi_{f}(\mathtt{i}\omega,\bm{k})$ at finite $\omega$ and $\bm{k}$, incorporating higher-order diagrams as done in Refs. [34, 46, 44], albeit these calculations were only carried out under the limit $\bm{k}\rightarrow 0$ for conductivity. Such an endeavor proves to be technically challenging, and we defer it to future studies.
The entanglement entropy in the CFL phase at $\nu=1/2$ has been numerically investigated in Ref. [94, 95]. Both of their results exhibit the $L\log(L)$ scaling, resembling that of free fermions, although their overall coefficients differ by a factor of two. This clearly deviates from our result of a boundary law $\mathcal{F}^{c}\sim L$ for bipartite charge fluctuations in Sec. IV.2.1. As we have seen in the gapless Mott insulator, although the charge fluctuations scale as $\mathcal{F}^{c}\sim L$, the spinon Fermi surface still gives rise to the spin fluctuations $\mathcal{F}^{s}\sim L\log(L)$. Therefore, it would be interesting to investigate the bipartite fluctuations of other quantities in the CFL phase that potentially identify the scaling of its entanglement entropy.
Although in this paper, we mainly discuss transitions at half-filling, the critical theories for the CFL-FL transition at other filling factors can be easily formulated within the vortex theory framework described in Sec. III. Specifically, the case of $\nu=1/4$ can be described by Eq. (19) together with the vortex sector
[…]
Through the fermionic particle-vortex duality [82, 83], this theory can be demonstrated to be dual to the critical theory proposed in Ref. [56] based on a parton construction, where the chargon is fractionalized into four fermions. At the critical point, we again observe the bipartite charge fluctuations described by Eq. (16), which exhibit a universal corner contribution related to the critical resistivity jumps via Eq. (15).
We have not considered the scenario of transitions between Landau FLs and charge-density-wave insulators coexisting with neutral Fermi surfaces [54, 56]. In such cases, the coupling of the density-wave order parameter $O$ to the Fermi surface of $f$-fermions leads to a different Landau damping term $\int_{k}|\omega||O(\omega,\mathbf{k})|^{2}$, which becomes relevant if the scaling dimension satisfies $\Delta[O]<1$. This scenario is likely true, and there is no dynamical decoupling of the boson and fermion sectors at criticality. It is possible that the Landau damping eventually drives the transition to become weakly first-order. Alternatively, there could be a new fixed point for the charge sector with a dynamical exponent $z>1$, as a technically similar theory has been considered in Ref. [96]. At this stage, we lack the theoretical tools to definitively determine the nature of these transitions, so the study of bipartite fluctuations across these exotic quantum critical points will need to be addressed in future research.
This paper focuses on critical Fermi surfaces in two spatial dimensions. However, continuous Mott transitions between Fermi-liquid metals and spin-liquid insulators in three-dimensional systems can be constructed using similar techniques. For instance, in Ref. [97], the same parton construction as described in Eq. (39) was employed, resulting in a critical theory similar to the two-dimensional case. In this context, the spinon-chargon interaction is found to be marginally irrelevant, and the chargon condensation belongs to the 4D XY universality class. Since the system reaches its upper critical dimension, the scaling forms of various physical quantities exhibit logarithmic corrections. Within the vortex theory framework, the U(1) disorder operator is represented by a Wilson surface. When examining its shape dependence, additional singular geometries arise, including corners, cones, and trihedral vertices. It would be intriguing to explore the universal terms in bipartite fluctuations for critical Fermi surfaces in three dimensions.
In Sec. III, we provide a unified theoretical framework for critical Fermi surfaces by introducing vortices in Landau FLs. A conceptually equivalent but technically distinct approach exists for constructing the vortex theory without using partons. To illustrate this idea, consider the CFL-FL transition. One can start with the non-linear bosonization of spinless FLs [47] (also see App. G for a brief review). Vortex excitations are then introduced in the following theory
$\mathcal{L}=\mathcal{L}_{\textrm{FS}}[\phi]+(a+A)\cdot J[\phi]-\frac{\mathtt{i}}{2\pi}\tilde{a}\wedge\textrm{d}a+\mathcal{L}_{\textrm{vrtx}}[\psi,\tilde{a}].$ (53)
Here, $\mathcal{L}_{\textrm{FS}}[\phi]$ is the nonlinear theory introduced in Eq. (118), and $J[\phi]$ represents the U(1) current. The dynamical gauge field $a_{\mu}$ is introduced as a Lagrangian multiplier, enforcing the duality relation $J[\phi]=\frac{\mathtt{i}}{2\pi}\star\textrm{d}\tilde{a}$. By considering the vortex sector as in Eq. (26), one can again realize the CFL-FL transition. The essence lies in the fact that the boson $\phi$ encompasses a large number of low-energy excitations, including modes represented by the partons. The advantage of the theory in Eq. (53) lies in its ability to undergo various generalizations by modifying $J[\phi]$, such as for higher-angular momentum channels, as noted in Ref. [98].
One of our key findings highlights the change in scaling behaviors of bipartite charge fluctuations (or the U(1) disorder operator) across interaction-driven “metal-insulator transitions,” which closely mirrors transitions between different phases of higher-form symmetries. Despite recent progress in understanding certain aspects of the low-energy physics of Fermi surfaces from the perspective of LU(1) anomaly [45], and efforts to formulate Fermi-surface dynamics using nonlinear bosonization via the infinite-dimensional Lie group of canonical transformations [47] (also see App. G), a generalized symmetry principle for the quantum phase transitions considered in this paper remains an open question. Nevertheless, our quantitative result Eq. (16) at the critical point, which identifies a universal corner contribution with a positive universal constant $C_{\rho}$, could be tested through quantum Monte Carlo simulations, potentially aiding in the discovery of UV lattice realizations. Given the recent numerical evidence of the CFL phase in twisted $\textrm{MoTe}_{2}$ [57, 58] and the CFL-FL transition reported in Ref. [71], it would be intriguing to explore the possibility of continuous CFL-FL transitions in related models. Numerically verifying Eq. (16) would provide strong evidence for continuous transitions and offer compelling support for the universal critical theory proposed in Refs. [55, 56].
Acknowledgment
We thank Luca Delacrétaz, Hart Goldman, Chao-Ming Jian, Prashant Kumar, Zlatko Papić, Dam Thanh Son, and Senthil Todadri for related discussions. We thank Cenke Xu for discussions and participation in the early stage of the project, and Meng Cheng for communicating unpublished results. This work was supported in part by the Simons Collaboration on Ultra-Quantum Matter, which is a grant from the Simons Foundation (651440), and the Simons Investigator award (990660).
Note added: (1) We would like to draw the reader’s attention to a related paper [99] that appeared in the same arXiv listing as the first version of this manuscript. (2) We would like to mention a subsequent work [88] that numerically verifies the predictions for CFLs (Eq. (12) along with Eq. (13)) through wavefunction-based Monte Carlo simulations. (3) We thank Kang-Le Cai, Meng Cheng, and Prashant Kumar for a related collaboration [75], which addresses the corner term in charge insulators and its relation to many-body quantum geometry.
Appendix A $\textrm{LU}(1)$ Anomaly and Structure Factor
In this appendix, we provide a geometric interpretation of the static structure factor, drawing from the $\textrm{LU}(1)$ anomaly of Fermi-surface states [45, 46]. We begin by introducing the low-energy patch theory
$\mathcal{L}=\int\textrm{d}\theta\psi^{\dagger}(\tau,\bm{x},\theta)(\textrm{D}_{\tau}+\mathtt{i}v_{F}^{j}(\theta)\textrm{D}_{j}+\kappa_{ij}(\theta)\textrm{D}_{i}\textrm{D}_{j})\psi(\tau,\bm{x},\theta)+(\textrm{interactions}).$ (54)
Here, $\theta$ denotes an angle variable labeling the patch, $v_{F}^{j}(\theta)$ represents the fermi velocity, and $\kappa_{ij}(\theta)$ is the curvature tensor. The gauge covariant derivative is denoted by $\textrm{D}_{\mu}=\partial_{\mu}-\mathtt{i}A_{\mu}$, where $A_{\mu}$ is the background electromagnetic field. Under the scaling limit, the paramagnetic current (at $A=0$) is given by
$J^{\tau}(\tau,\bm{x})$ $=\frac{\delta\mathcal{S}}{\delta A_{\tau}}=-\mathtt{i}\int\textrm{d}\theta\rho(\tau,\bm{x},\theta),$ (55)
$J^{i}(\tau,\bm{x})$ $=\frac{\delta\mathcal{S}}{\delta A_{i}}=\int\textrm{d}\theta\rho(\tau,\bm{x},\theta)v_{F}^{i}(\theta),$ (56)
where $\rho(\tau,\bm{x},\theta)=\psi^{\dagger}(\tau,\bm{x},\theta)\psi(\tau,\bm{x},\theta)$ represents the density at each patch of the Fermi surface. We introduce the phase-space current density $\mathcal{J}^{\mu}$ such that $J^{\mu}(\tau,\bm{x})=\int\textrm{d}\theta\mathcal{J}^{\mu}(\tau,\bm{x},\theta)$. Due to the $\textrm{LU}(1)$ anomaly [45], the Ward identity (in Euclidean signature) is as follows
$\partial_{\mathtt{I}}\mathcal{J}^{\mathtt{I}}=\frac{-\mathtt{i}}{8\pi^{2}}\varepsilon^{\mathtt{IJKL}}\partial_{\mathtt{I}}\mathcal{A}_{\mathtt{J}}\partial_{\mathtt{K}}\mathcal{A}_{\mathtt{L}},$ (57)
where $\partial_{\mathtt{I}}=(\partial_{\tau},\partial_{\bm{x}},\partial_{\theta})$. As for the phase-space background field $\mathcal{A}_{\mathtt{I}}$, we introduce
$\mathcal{A}_{\tau}=A_{\tau}(\tau,\bm{x}),\qquad\mathcal{A}_{\bm{x}}=A_{\bm{x}}(\tau,\bm{x})+\bm{k}_{F}(\theta),\qquad\mathcal{A}_{\theta}\textrm{ is independent of }\tau,\bm{x}$ (58)
Here, $A_{\mu}=(A_{\tau},A_{\bm{x}})$ is the ordinary background electromagnetic field. We provide two explanations for the inclusion of $\bm{k}_{F}(\theta)$ in $\mathcal{A}_{\bm{x}}$. (1) We want to turn on a background flux $F_{\bm{x}\bm{k}}$ that ensures the canonical commutation relation $[x_{i},k_{j}]=\mathtt{i}\delta_{ij}$ [100]. (2) It aligns with the semiclassical equation of motion $\dot{\bm{k}}=E$. (One may also check Sec. VI. B of Ref. [45]). The LU(1) Ward identity Eq. (57) leads to
$\partial_{\tau}\mathcal{J}^{\tau}(\tau,\bm{x},\theta)+\partial_{\bm{x}}\cdot\mathcal{J}^{\bm{x}}(\tau,\bm{x},\theta)=\frac{-\mathtt{i}}{(2\pi)^{2}}\bm{E}\times\frac{\textrm{d}\bm{k}_{F}(\theta)}{\textrm{d}\theta},$ (59)
where $\bm{E}=-F_{\tau\bm{x}}=\partial_{\bm{x}}A_{\tau}-\partial_{\tau}A_{\bm{x}}$. We have used $\partial_{\theta}J^{\theta}=0$ in the presence of a background electric field [45]. We focus on the response to $A_{\tau}$ and set $A_{\bm{x}}=0$. In the momentum space, this is given by
$(-\mathtt{i}\omega+\bm{v}_{F}\cdot\bm{k})\rho(\omega,\bm{k},\theta)=\frac{-\mathtt{i}}{(2\pi)^{2}}\left(\bm{k}\times\frac{\textrm{d}\bm{k}_{F}(\theta)}{\textrm{d}\theta}\right)A_{\tau}(\omega,\bm{k}).$ (60)
The equal-time response of the total electric charge density $J^{\tau}(\omega,\bm{k})=-\mathtt{i}\int\textrm{d}\theta\rho(\omega,\bm{k},\theta)$ is therefore
$\Pi^{\tau\tau}(\tau\rightarrow 0,\bm{k})=\frac{1}{(2\pi)^{2}}\int\frac{\textrm{d}\omega}{2\pi}\int_{0}^{2\pi}\textrm{d}\theta\frac{1}{\mathtt{i}\omega-\bm{v}_{F}(\theta)\cdot\bm{k}}\left(\bm{k}\times\frac{\textrm{d}\bm{k}_{F}(\theta)}{\textrm{d}\theta}\right)=-\frac{\textrm{Area}}{(2\pi)^{2}},$ (61)
[…]
For a spherical Fermi surface $\bm{k}_{F}(\theta)=k_{F}(\cos\theta,\sin\theta)$, we have $\bm{k}\times\partial_{\theta}\bm{k}_{F}(\theta)=\bm{k}\cdot\bm{k}_{F}(\theta)$, and accordingly
$\Pi^{\tau\tau}(\tau\rightarrow 0,\bm{k})=\frac{1}{(2\pi)^{2}}\int\frac{\textrm{d}\omega}{2\pi}\int_{0}^{2\pi}\textrm{d}\theta\frac{k_{F}|\bm{k}|\cos\theta}{\mathtt{i}\omega-v_{F}|\bm{k}|\cos\theta}=\frac{k_{F}|\bm{k}|}{2\pi}\int_{0}^{2\pi}\frac{\textrm{d}\theta}{2\pi}\left(-\frac{|\cos\theta|}{2}\right)=-\frac{k_{F}|\bm{k}|}{2\pi^{2}}.$ (62)
This exactly reproduces the free-fermion result in Eq. (3). It’s worth noting that one can introduce interactions that explicitly break the LU(1) symmetry, such as scattering between fermion modes from different patches. For instance, considering the RPA treatment of a local density-density interaction, we obtain the density response
$\hat{\Pi}^{\tau\tau}(\mathtt{i}\omega,\bm{k})=\frac{\Pi^{\tau\tau}(\mathtt{i}\omega,\bm{k})}{1-u\Pi^{\tau\tau}(\mathtt{i}\omega,\bm{k})},$ (63)
Here, $\Pi^{\tau\tau}(\mathtt{i}\omega,\bm{k})$ represents the free-fermion result from Eq. (30), and $u$ denotes a coupling constant. Consequently, the equal-time function becomes
$\int\frac{\textrm{d}\omega}{2\pi}\hat{\Pi}^{\tau\tau}(\mathtt{i}\omega,\bm{k})=-\frac{k_{F}|\bm{k}|}{2\pi^{2}}\frac{\sqrt{2\tilde{u}+1}+2\tilde{u}\arctan(\sqrt{2\tilde{u}+1})}{(2\tilde{u}+1)^{3/2}},$ (64)
where $\tilde{u}=u\mathscr{D}_{F}$ is rescaled by the density of states $\mathscr{D}_{F}=\frac{k_{F}}{2\pi v_{F}}$ at the fermi level. Also, note that in the problem of the half-filled Landau level, the Ward identity (Eq. (57)) needs modification due to the presence of an emergent gauge field [45]. Therefore, the geometric interpretation presented here should not conflict with the results in Sec. IV.2.1.
Appendix B Coulomb Gauge and Response Theory
In this appendix, we introduce our convention about the response theory under the Coulomb gauge. For any rotationally invariant systems, the two-point function $\Pi^{\mu\nu}=\langle J^{\mu}J^{\nu}\rangle$ of conserved currents has the structure
$\Pi^{\tau\tau}(k)$ $=\frac{|\bm{k}|^{2}}{\omega^{2}}\Pi_{L}(\omega,|\bm{k}|),$
$\Pi^{\tau i}(k)$ $=-\frac{k_{i}}{\omega}\Pi_{L}(\omega,|\bm{k}|)+\varepsilon^{ij}k_{j}\mathsf{H}(\omega,|\bm{k}|),$
$\Pi^{ij}(k)$ $=\frac{k_{i}k_{j}}{|\bm{k}|^{2}}\Pi_{L}(\omega,|\bm{k}|)+\left(\delta^{ij}-\frac{k_{i}k_{j}}{|\bm{k}|^{2}}\right)\Pi_{T}(\omega,|\bm{k}|)+\omega\varepsilon^{ij}\mathsf{H}(\omega,|\bm{k}|),$ (65)
where $\Pi_{L}$ and $\Pi_{T}$ represent the longitudinal and transverse components respectively, while $\mathsf{H}$ governs the Hall response. It is sometimes beneficial to decompose the spatial components of the background/gauge field into longitudinal and transverse components, expressed as $\bm{A}=\bm{A}_{L}+\bm{A}_{T}$ where
$\bm{A}^{L}$ $(k)=\mathcal{P}^{L}\bm{A}(k),\qquad\mathcal{P}_{ij}^{L}=\frac{k_{i}k_{j}}{|\bm{k}|^{2}},$
$\bm{A}^{T}$ $(k)=\mathcal{P}^{T}\bm{A}(k),\qquad\mathcal{P}_{ij}^{T}=\delta_{ij}-\mathcal{P}_{ij}^{L}.$ (66)
It is useful to introduce a scalar $A_{T}$ to represent the transverse fields $A_{i}^{T}$, which is defined by
$A_{i}^{T}(k)=\frac{\varepsilon_{ij}k_{j}}{|\bm{k}|}A_{T}(k),\qquad A_{T}(k)=\frac{\varepsilon_{ij}k_{j}}{|\bm{k}|}A_{i}^{T}(k).$ (67)
Under the Coulomb gauge $\bm{A}^{L}=0$, the response theory can expressed in the basis $(A_{\tau},A_{T})$
$\mathcal{S}[A]=\int\frac{\textrm{d}^{3}k}{(2\pi)^{3}}\frac{-1}{2}(A_{\tau}(-k)\;A_{T}(-k))\left(\begin{array}[]{cc}\Pi^{\tau\tau}(k)&\mathsf{H}(k)|\bm{k}|\\ \mathsf{H}(k)|\bm{k}|&\Pi^{TT}(k)\end{array}\right)\left(\begin{array}[]{c}A_{\tau}(k)\\ A_{T}(k)\end{array}\right),$ (72)
where $\Pi^{\tau\tau}(k)=\frac{|\bm{k}|^{2}}{\omega^{2}}\Pi_{L}(\omega,|\bm{k}|)$ and $\Pi^{TT}(k)=-\Pi_{T}(k)$. Here, $k=(\omega,\bm{k})$ is a collective notion of frequency and momentum. Notice that the Chern-Simons term is
$\int\frac{-\mathtt{i}}{4\pi}A\wedge\textrm{d}A=\int\frac{\textrm{d}^{3}k}{(2\pi)^{3}}\frac{-1}{2}(A_{\tau}(-k)\;A_{T}(-k))\frac{|\bm{k}|}{2\pi}\left(\begin{array}[]{cc}0&1\\ 1&0\end{array}\right)\left(\begin{array}[]{c}A_{\tau}(k)\\ A_{T}(k)\end{array}\right).$ (77)
Appendix C Ioffe-Larkin Composition Rule
We investigate the response function of electrons using the parton theory described in Eq. (18). In the standard RPA approach, integrating out both fermionic and bosonic partons $f$ and $b$ leads to
$\mathcal{S}=\int_{k}\frac{-\Pi_{f}^{\mu\nu}(k)}{2}(a_{\mu}(-k)+e_{f}A_{\mu}(-k))(a_{\nu}(k)+e_{f}A_{\nu}(k))+\frac{-\Pi_{b}^{\mu\nu}(k)}{2}(-a_{\mu}(-k)+e_{b}A_{\mu}(-k))(-a_{\nu}(k)+e_{b}A_{\nu}(k)).$ (78)
To enforce gauge invariance, we integrate out the dynamical gauge fields $a_{\mu}$ at the RPA level. This yields the total response theory $\mathcal{L}=\int_{k}\frac{-1}{2}\Pi^{\mu\nu}(k)A_{\mu}(-k)A_{\nu}(k)$ with
$\Pi$ $=e_{f}^{2}\Pi_{f}+e_{b}^{2}\Pi_{b}-(e_{f}\Pi_{f}-e_{b}\Pi_{b})(\Pi_{f}+\Pi_{b})^{-1}(e_{f}\Pi_{f}-e_{b}\Pi_{b})$
$=e_{f}^{2}(\Pi_{f}-\Pi_{f}(\Pi_{f}+\Pi_{b})^{-1}\Pi_{f})+e_{b}^{2}(\Pi_{b}-(\Pi_{f}+\Pi_{b})^{-1}\Pi_{b})+e_{b}e_{f}(\Pi_{f}(\Pi_{f}+\Pi_{b})^{-1}\Pi_{b}+\Pi_{b}(\Pi_{f}+\Pi_{b})^{-1}\Pi_{f})$
$=(e_{f}+e_{b})^{2}\Pi_{f}(\Pi_{f}+\Pi_{b})^{-1}\Pi_{b}=\Pi_{f}(\Pi_{f}+\Pi_{b})^{-1}\Pi_{b},$ (79)
where we have used $e_{f}+e_{b}=1$. Therefore, we find the Ioffe-Larkin composition rule
$\Pi^{-1}=\Pi_{f}^{-1}+\Pi_{b}^{-1},$ (80)
which is independent of the assignment $(e_{f},e_{b})$ of the global $\textrm{U}(1)$ charge.
Under the Coulomb gauge, the parton response functions $\Pi_{f}$ and $\Pi_{b}$ can be written as
$\Pi_{f}=\left(\begin{array}[]{cc}\Pi_{f}^{\tau\tau}(k)&\mathsf{H}_{f}|\bm{k}|\\ \mathsf{H}_{f}|\bm{k}|&\Pi_{f}^{TT}(k)\end{array}\right),\qquad\Pi_{b}=\left(\begin{array}[]{cc}\Pi_{b}^{\tau\tau}(k)&\mathsf{H}_{b}|\bm{k}|\\ \mathsf{H}_{b}|\bm{k}|&\Pi_{b}^{TT}(k)\end{array}\right).$ (81)
The Ioffe-Larkin rule Eq. (80) leads to the gauge-invariant response
$\Pi^{\tau\tau}$ $=\frac{\Pi_{f}^{\tau\tau}\det(\Pi_{b})+\Pi_{b}^{\tau\tau}\det(\Pi_{f})}{\det(\Pi_{f}+\Pi_{b})},$
$\Pi^{TT}$ $=\frac{\Pi_{f}^{TT}\det(\Pi_{b})+\Pi_{b}^{TT}\det(\Pi_{f})}{\det(\Pi_{f}+\Pi_{b})},$
$\Pi^{\tau T}$ $=\Pi^{T\tau}=|\bm{k}|\frac{\mathsf{H}_{f}\det(\Pi_{b})+\mathsf{H}_{b}\det(\Pi_{f})}{\det(\Pi_{f}+\Pi_{b})}.$ (82)
For time-reversal invariant systems such that $\mathsf{H}_{f}=0$ and $\mathsf{H}_{b}=0$, one simply has
$\Pi^{\tau\tau}=\frac{\Pi_{f}^{\tau\tau}\Pi_{b}^{\tau\tau}}{\Pi_{f}^{\tau\tau}+\Pi_{b}^{\tau\tau}},\qquad\Pi^{TT}=\frac{\Pi_{f}^{TT}\Pi_{b}^{TT}}{\Pi_{f}^{TT}+\Pi_{b}^{TT}},\qquad\Pi^{\tau T}=\Pi^{T\tau}=0.$ (83)
Appendix D Scaling of Bipartite Fluctuations
In this appendix, we present the technical details of gauge-invariant regularization schemes used to handle the UV divergence in evaluating Eq. (25). Our starting point is the static structure factor of electrons. At long wavelengths, we assume rotational invariance and consider a generic power-law correlation Eq. (2) in real space. To keep our discussion as general as possible, we allow the value of $\alpha$ to be continuously tuned with $\alpha>2$. It is known that the overall coefficient $C_{0}$ is universal in electronic systems described by CFTs where $\alpha=4$, while it is sensitive to interactions in Landau FLs where $\alpha=3$. In App. D.1, we propose a real-space method that is suitable for both dimensional regularization and UV cut-off schemes.
After the Fourier transformation of Eq. (2), the static structure factor is given by
$\Pi^{\tau\tau}(\tau\rightarrow 0,\bm{k})=\int\textrm{d}^{d}x\frac{C_{0}}{|\bm{x}|^{\alpha}}e^{\mathtt{i}\bm{k}\cdot\bm{x}}=\frac{2^{d-\alpha}\pi^{\frac{d}{2}}\Gamma(\frac{d-\alpha}{2})}{\Gamma(\frac{\alpha}{2})}\frac{C_{0}}{\left|\bm{k}\right|^{d-\alpha}}=\begin{cases}+2\pi C_{0}\log(1/|\bm{k}|)&\alpha=2\\ -2\pi C_{0}|\bm{k}|&\alpha=3\\ -\frac{\pi}{2}C_{0}|\bm{k}|^{2}\log(1/|\bm{k}|)&\alpha=4\\ +\frac{2\pi}{9}C_{0}|\bm{k}|^{3}&\alpha=5\\ +\frac{\pi}{32}C_{0}|\bm{k}|^{4}\log(1/|\bm{k}|)&\alpha=6\\ \qquad\vdots&\quad\vdots\end{cases}$ (84)
where $d=2$. In the case of gapless modes in CFLs, there is a contribution of the form $|\bm{k}|^{3}\log(1/|\bm{k}|)$, which results in an additional term $|\bm{x}|^{-5}\log|\bm{x}|$ alongside Eq. (2) with $\alpha=5$. However, this logarithmic correction does not change the scaling of bipartite fluctuations. This issue is further discussed in App. D.2.
D.1 Real-Space Method
We embed the equal-time density response Eq. (2) into a space-time current-current correlation
$\langle\check{J}^{\mu}(x)\check{J}^{\nu}(0)\rangle=\frac{C_{0}}{|x|^{\alpha}}\left(\delta^{\mu\nu}-\frac{\alpha}{\alpha-2}\frac{x^{\mu}x^{\nu}}{|x|^{2}}\right),$ (85)
such that $\Pi^{\tau\tau}(\tau\rightarrow 0,\bm{x})=\langle\check{J}^{\tau}(\tau,\bm{x})\check{J}^{\tau}(\tau,0)\rangle$. We once again introduce a dual gauge field $\check{a}_{\mu}$ to represent the current $\check{J}^{\mu}=\frac{\mathtt{i}}{2\pi}\varepsilon^{\mu\nu\rho}\partial_{\nu}\check{a}_{\rho}$. The gauge-field propagator $\check{D}_{\mu\nu}(x-y)=\langle\check{a}_{\mu}(x)\check{a}_{\nu}(y)\rangle$ can be written as
$\check{D}_{\mu\nu}^{(\alpha)}(x)$ $=\frac{(2\pi)^{2}C_{0}}{(\alpha-2)^{2}}\frac{1}{|x|^{\alpha-2}}\left((1+\zeta)\delta^{\mu\nu}-\zeta(\alpha-2)\frac{x^{\mu}x^{\nu}}{|x|^{2}}\right)$ (86)
$=\frac{(2\pi)^{2}C_{0}}{(\alpha-2)^{2}}\left(\frac{\delta^{\mu\nu}}{|x|^{\alpha-2}}+\zeta\partial_{\mu}\partial_{\nu}\begin{cases}\frac{1}{4-\alpha}|x|^{4-\alpha}&\alpha\neq 4\\ \log|x|&\alpha=4\end{cases}\right),$ (87)
where $\zeta$ is a Faddeev-Popov gauge-fixing parameter. One can replace the gauge field $\tilde{a}_{\mu}$ by $\check{a}_{\mu}$ in the Wilson loop Eq. (24), and calculate the bipartite fluctuations Eq. (25) by
$\mathcal{F}_{\mathcal{A}}^{(\alpha,\epsilon)}=\frac{1}{(2\pi)^{2}}\int_{\mathcal{C}}\textrm{d}x^{i}\int_{\mathcal{C}}\textrm{d}y^{j}\check{D}_{ij}^{(\alpha)}(\epsilon,\bm{x}-\bm{y}),$ (88)
where a small splitting $\epsilon>0$ in the “temporal direction” serves as a small real-space UV cut-off, and the integrals are performed along the closed spatial loop $\mathcal{C}=\partial\mathcal{A}$. In Eq. (88), there are two parameters, $\epsilon$ and $\alpha$. If the cut-off $\epsilon$ is strictly set to zero, and the power $\alpha$ matches the physical value from the Ioffe-Larkin rule Eq. (23), the generalized formula Eq. (88) exactly reduces back to the original definition Eq. (25).
Now, we are prepared to address the gauge-invariant calculation of Eq. (88). In the UV cut-off scheme, we fix the value of $\alpha$ and treat $\epsilon$ as a small expansion parameter. Let us examine the square geometry of the loop $\mathcal{C}$, which has a side length of $L$. The case of CFTs where $\alpha=4$ has been extensively discussed in Ref. [18]. It has a leading cut-off-dependent boundary-law term together with a universal subleading logarithmic term
$\frac{\mathcal{F}_{\mathcal{A}}^{(4,\epsilon)}}{C_{0}}=\frac{\pi}{4}\frac{|\mathcal{C}|}{\epsilon}-2\log|\mathcal{C}|+\textrm{const},$ (89)
where $|\mathcal{C}|=4L$ is the perimeter of the square. By setting $\alpha=3$ and following the calculations in Eq. (9)-(12) from Ref. [18], we obtain the result for Fermi liquids
$\frac{\mathcal{F}_{\mathcal{A}}^{(3,\epsilon)}}{C_{0}}=2|\mathcal{C}|(\log|\mathcal{C}|+\textrm{const}),$ (90)
As a self-consistency check, the final result of $\mathcal{F}_{\mathcal{A}}$ is again independent of the Faddeev-Popov gauge-fixing parameter $\zeta$. We have also evaluated Eq. (88) for generic values of $\alpha$ other than 3 and 4
$\frac{\mathcal{F}_{\mathcal{A}}^{(\alpha,\epsilon)}}{C_{0}}=|\mathcal{C}|\epsilon^{3-\alpha}\frac{\sqrt{\pi}\Gamma(\frac{\alpha-3}{2})}{(\alpha-2)^{2}\Gamma(\frac{\alpha-2}{2})}+|\mathcal{C}|^{4-\alpha}\frac{(2^{\frac{\alpha}{2}}-2-2(\alpha-4)\textrm{Hypergeometric2F1}(1,\frac{3-\alpha}{2};\frac{3}{2};-1))}{2^{5-\frac{3\alpha}{2}}\Gamma(\alpha-1)/\Gamma(\alpha-4)}.$ (91)
The general expression Eq. (91) is informative in understanding the contribution of gapless modes in different phases of matter. In non-local systems where the instantaneous charge correlation decays even slower than Landau Fermi liquids (i.e., $\alpha<3$), the first boundary-law term in Eq. (91) vanishes as $\epsilon$ approaches zero, leaving only the second term, which is independent of $\epsilon$ and scales as $|\mathcal{C}|^{4-\alpha}$. Conversely, when $\alpha>3$, the leading-order term is always given by the boundary law $|\mathcal{C}|\epsilon^{3-\alpha}$, which contains a power-law UV divergence. Another notable observation is that when the instantaneous charge correlation is weaker than in CFTs (i.e., $\alpha>4$), the universal subheading term $|\mathcal{C}|^{4-\alpha}$ vanishes in the large-$|\mathcal{C}|$ limit. As we will see, this holds true for CFLs. In Tab. 1, we summarize our findings under the cut-off regularization scheme.
If one is only interested in the “universal term” that remains independent of any UV cut-off, there is another convenient regularization scheme that is in the same spirit as dimensional regularization. Here, we set $\epsilon=0$ and retain $\alpha$ as an arbitrary parameter in the integrals for Eq. (88). Subsequently, we consider the final result through an expansion in terms of small $\delta=\alpha-\bar{\alpha}$, where $\bar{\alpha}$ represents the physical value. In this scheme, all power-law UV divergences are automatically eliminated, and the logarithmic divergence manifests as $\delta^{-1}$. Let us check two simple geometries, a square and a circle. One can easily find the result for a square
[…]
and the result for a circle
$\frac{\mathcal{F}_{\mathcal{A}}^{(\alpha,0)}}{C_{0}}$ $=|\mathcal{C}|^{4-\mathit{\alpha}}\frac{\pi^{\mathit{\alpha}-\frac{5}{2}}\Gamma(\frac{3}{2}-\frac{\mathit{\alpha}}{2})}{2(\alpha-2)\Gamma(3-\frac{\alpha}{2})}=\begin{cases}2|\mathcal{C}|(\log|\mathcal{C}|-\delta^{-1}-\log(\frac{\pi}{2}))&\alpha=3\\ -\frac{\pi^{2}}{2}&\alpha=4\\ -\frac{\pi^{2}}{3}|\mathcal{C}|^{-1}(\log|\mathcal{C}|-\delta^{-1}+\frac{1}{6}(5-\log(\frac{\pi^{6}}{64})))\rightarrow 0&\alpha=5\end{cases},$ (93)
where $|\mathcal{C}|$ represents the perimeter of $\mathcal{C}=\partial\mathcal{A}$ in the both cases. There are some lessons we can learn from this exercise. (1) When $\alpha=3$, the coefficient of the leading term $|\mathcal{C}|\log|\mathcal{C}|$ remains independent of detailed geometries. (2) Both the subleading term $|\mathcal{C}|$ in the case of $\alpha=3$ and the universal constant term in the case of $\alpha=4$ depend on detailed geometries. (3) When $\alpha=5$, the universal term depends on the geometry but always vanishes under the large-$|\mathcal{C}|$ scaling. (4) When $\alpha=4$, the term $\log|\mathcal{C}|$ only appears when the geometry contains sharp corners.
In fact, the corner contribution can also be conveniently calculated using this regularization scheme, compared to the calculation under the cut-off scheme in Ref. [18]. It is convenient to choose the gauge $\zeta=\frac{1}{\alpha-3}$, ensuring that contribution from the same straight line vanishes. Considering the correlation between two straight lines, we find that the contribution exhibits a logarithmic divergence only when $\alpha=3$ and $\alpha=4$. For $\alpha=4$, the angle dependence Eq. (6) can also be exactly reproduced.
D.2 Momentum-Space Method
Starting with the propagator $\tilde{D}_{\mu\nu}=\langle\tilde{a}_{\mu}\tilde{a}_{\nu}\rangle$ for the gauge field $\tilde{a}_{\mu}$ as defined in Eq. (20), the bipartite fluctuations described in Eq. (25) can be rewritten as
$\mathcal{F}_{\mathcal{A}}=\frac{1}{(2\pi)^{2}}\int\frac{\textrm{d}\omega\textrm{d}^{2}\bm{k}}{(2\pi)^{3}}Y_{\mathcal{C}}^{i}(-\bm{k})D_{ij}(\mathtt{i}\omega,\bm{k})Y_{\mathcal{C}}^{j}(\bm{k})$ (94)
by defining the linear functional over the closed loop $\mathcal{C}=\partial\mathcal{A}$ as
$Y_{\mathcal{C}}^{i}(\bm{k})=\int_{\mathcal{C}}\textrm{d}x^{i}e^{\mathtt{i}\bm{k}\cdot\bm{x}}.$ (95)
Taking advantage of the divergence-free property $k_{i}Y_{\mathcal{C}}^{i}(\bm{k})=-\mathtt{i}\int_{\mathcal{C}}\textrm{d}x^{i}\partial_{i}e^{\mathtt{i}\bm{k}\cdot\bm{x}}=0$, we can write $Y_{\mathcal{C}}^{i}=\varepsilon^{ij}k_{j}\overline{Y}_{\mathcal{C}}$, where
$\overline{Y}_{\mathcal{C}}(\bm{k})=\mathtt{i}\int_{\mathcal{A}}\textrm{d}^{2}\bm{x}e^{\mathtt{i}\bm{k}\cdot\bm{x}}.$ (96)
Consequently, Eq. (94) can be equivalently expressed as
$\mathcal{F}_{\mathcal{A}}=\int\frac{\textrm{d}^{2}\bm{k}}{(2\pi)^{2}}\Pi^{\tau\tau}(\tau\rightarrow 0,\bm{k})|\overline{Y}_{\mathcal{C}}(\bm{k})|^{2},$ (97)
where $\Pi^{\tau\tau}(\tau\rightarrow 0,\bm{k})$ is the static structure factor.
We are particularly interested in the term $\left|\bm{k}\right|^{\beta}$ or $\left|\bm{k}\right|^{\beta}\log|\bm{k}|$ that appears in the static structure factor. We find that the Fourier transform of $\left|\bm{k}\right|^{\beta}$ is nonzero only when $\beta$ is an odd integer
$\int\frac{\textrm{d}^{d}\bm{k}}{(2\pi)^{d}}e^{-\mathtt{i}\bm{k}\cdot\bm{x}}\left|\bm{k}\right|^{\beta}=\frac{2^{\beta}\pi^{\frac{1-d}{2}}\Gamma(\frac{d+\beta}{2})}{\Gamma(\frac{d-1}{2})\Gamma(1-\frac{d+\beta}{2})}\frac{1}{|\bm{x}|^{d+\beta}}=\begin{cases}0&\beta=0\\ -\frac{1}{2\pi}|\bm{x}|^{-3}&\beta=1\\ 0&\beta=2\\ +\frac{9}{2\pi}|\bm{x}|^{-5}&\beta=3\\ 0&\beta=4\\ \qquad\vdots&\quad\vdots\end{cases}$ (98)
where $d=2$. As for $\left|\bm{k}\right|^{\beta}\log|\bm{k}|$, we find nonzero results for all integer values of $\beta$
$\int\frac{\textrm{d}^{d}\bm{k}}{(2\pi)^{d}}e^{-\mathtt{i}\bm{k}\cdot\bm{x}}\left|\bm{k}\right|^{\beta}\log|\bm{k}|$ $=\frac{2^{\beta}\pi^{\frac{1-d}{2}}\Gamma(\frac{d+\beta}{2})}{\Gamma(\frac{d-1}{2})\Gamma(1-\frac{d+\beta}{2})}\frac{\Psi^{(0)}(\frac{d+\beta}{2})+\Psi^{(0)}(1-\frac{d+\beta}{2})-2\log(|\bm{x}|/2)}{2|\bm{x}|^{d+\beta}}$
=$=\begin{cases}-\frac{1}{2\pi|\bm{x}|^{2}}&\beta=0\\ +\frac{\log|\bm{x}|+\log 2+\gamma_{E}-2}{2\pi|\bm{x}|^{3}}&\beta=1\\ +\frac{2}{\pi|\bm{x}|^{4}}&\beta=2\\ -\frac{3(3\log|\bm{x}|+\log 8+3\gamma_{E}-8)}{2\pi|\bm{x}|^{5}}&\beta=3\\ -\frac{32}{\pi|\bm{x}|^{6}}&\beta=4\\ \qquad\vdots&\quad\vdots\end{cases}$ (99)
where again $d=2$, $\Psi^{(0)}(z)=\frac{\textrm{d}}{\textrm{d}z}\log\Gamma(z)$ denotes the the polygamma function, and $\gamma_{E}$ is the Euler–Mascheroni constant. The cases from Eq. (98) with odd $\beta$ and those from Eq. (99) with even $\beta$ together account for all the cases considered earlier in Eq. (84). The new cases arise in Eq. (99) with odd $\beta$, where the power-law correlation receives a logarithmic correction.
We can evaluate Eq. (97) using dimensional regularization, treating $\beta$ as a tunable parameter in the integral. To illustrate this, consider a circle $\mathcal{C}$ with a radius $R$. In this case, we have
$\overline{Y}_{\mathcal{C}}(\bm{k})=\mathtt{i}\int_{0}^{R}r\textrm{d}r\int_{0}^{2\pi}\textrm{d}\theta e^{\mathtt{i}|\bm{k}|r\cos\theta}=\frac{\mathtt{i}2\pi R}{|\bm{k}|}\textrm{BesselJ}(1,|\bm{k}|R).$ (100)
For the term proportional to $\left|\bm{k}\right|^{\beta}$, the contribution to to bipartite fluctuations is given by
$\mathcal{F}_{\mathcal{A}}\sim\int\frac{\textrm{d}^{2}\bm{k}}{(2\pi)^{2}}|\bm{k}|^{\beta}|\overline{Y}_{\mathcal{C}}(\bm{k})|^{2}=2\pi R^{2}\int_{0}^{+\infty}\textrm{d}kk^{\beta-1}\textrm{BesselJ}(1,|\bm{k}|R)^{2}=\frac{\sqrt{\pi}\Gamma(1+\frac{\beta}{2})\Gamma(\frac{1}{2}-\frac{\beta}{2})}{\Gamma(1-\frac{\beta}{2})\Gamma(2-\frac{\beta}{2})}R^{2-\beta}.$ (101)
For the term proportional to $\left|\bm{k}\right|^{\beta}\log|\bm{k}|$, the contribution becomes
$\mathcal{F}_{\mathcal{A}}$ $\sim\int\frac{\textrm{d}^{2}\bm{k}}{(2\pi)^{2}}|\bm{k}|^{\beta}\log|\bm{k}||\overline{Y}_{\mathcal{C}}(\bm{k})|^{2}=2\pi R^{2}\int_{0}^{+\infty}\textrm{d}kk^{\beta-1}\log(k)\textrm{BesselJ}(1,|\bm{k}|R)^{2}$
$=\frac{\sqrt{\pi}\Gamma(1+\frac{\beta}{2})\Gamma(\frac{1}{2}-\frac{\beta}{2})(\frac{2}{\beta-2}+\Psi^{(0)}(\frac{1}{2}-\frac{\beta}{2})-2\Psi^{(0)}(1-\frac{\beta}{2})-\Psi^{(0)}(1+\frac{\beta}{2})+2\log R)}{(\beta-2)\Gamma(1-\frac{\beta}{2})^{2}}R^{2-\beta}.$ (102)
Applying these results to some examples of compressible states, we find the cut-off-independent contributions
$\begin{cases}|\bm{k}|\quad\Rightarrow\quad 2R\log R&\textrm{Landau FLs and superfluids}\\ |\bm{k}|^{3}\log(1/|\bm{k}|)\quad\Rightarrow\quad\frac{3}{8}\frac{(\log R)^{2}}{R}\rightarrow 0&\textrm{gapless modes in CFLs}\\ |\bm{k}|\log(1/|\bm{k}|)\quad\Rightarrow\quad R(\log R)^{2}&\textrm{certain non-Fermi liquids}\end{cases}.$ (103)
Therefore, in CFLs, the subleading term from gapless modes vanishes under large-scale scaling. For the leading term, since $|\bm{x}|^{-4}<|\bm{x}|^{-5}\log|\bm{x}|<|\bm{x}|^{-5}$ at long distances, and both $|\bm{x}|^{-4}$ and $|\bm{x}|^{-5}$ lead to a boundary law, the logarithmic correction does not change the leading-order scaling. This conclusion has also been independently confirmed in Ref. [99] using a momentum-space cut-off scheme. The correction involving $|\bm{x}|^{-3}\log|\bm{x}|$ (or $|\bm{k}|\log(1/|\bm{k}|)$) deserves special attention, as $|\bm{x}|^{-3}$ (or $|\bm{k}|$) already exhibits a logarithmic violation of the boundary law. An example of such “double-log violation” of the boundary law, $R(\log R)^{2}$, was first pointed out in Ref. [99].
Appendix E Dual Theories for CFL-FL Transition
We begin by establishing the duality relation between the vortex theory Eq. (19) (together with Eq. (26)) and the critical theory Eq. (29). It is known that a single Dirac fermion enjoys the fermion-fermion duality [83, 82]
[…]
where each Dirac fermion is defined through the Pauli-Villars scheme with another heavy Dirac fermion in the UV. Upon integrating out the gauge field $b$, the resulting expression takes on the usual form found in the literature, albeit with incorrectly quantized topological terms
[…]
Its time-reversal image yields yet another fermionic particle-vortex duality
[…]
Let us apply the duality Eq. (106) to the two fermions in Eq. (26), subject to the constraint $\psi_{1}^{\dagger}\psi_{1}=\psi_{2}^{\dagger}\psi_{2}$
[…]
where the fluxes of $\tilde{b}$ and $\tilde{c}$ represent the densities of $\psi_{1}$ and $\psi_{2}$, and $\lambda$ serves as a Lagrangian multiplier. After integrating out both $\lambda$ and $\tilde{a}$, we find that (when $\psi_{1}^{\dagger}\psi_{1}=\psi_{2}^{\dagger}\psi_{2}$)
[…]
Together with the Fermi-surface sector described by $\mathcal{L}_{\textrm{FS}}[f,a]$, we find Eq. (19) (together with Eq. (26)) and Eq. (29) are indeed related by the fermionic particle-vortex duality [83, 82].
In view of the abelian duality web [83], there are other formulations of the critical theory as well. For a single Dirac fermion, there is a fermion-boson particle-vortex duality
[…]
$|\textrm{D}_{\tilde{b}}\varphi_{1}|^{2}+|\varphi_{1}|^{4}+|\textrm{D}_{\tilde{c}}\varphi_{2}|^{2}+|\varphi_{2}|^{4}-\frac{\mathtt{i}}{4\pi}\tilde{b}\textrm{d}\tilde{b}-\frac{\mathtt{i}}{4\pi}\tilde{c}\textrm{d}\tilde{c}-\frac{\mathtt{i}}{2\pi}\tilde{a}\textrm{d}(a+\tilde{b}+\tilde{c}).$ (110)
Integrating out $\tilde{a}$ imposes the constraint $\tilde{c}=-a-\tilde{b}$, leading to the bosonic dual theory for the CFL-FL transition
$\mathcal{L}=\mathcal{L}_{\textrm{FS}}[f,a+A]+|\textrm{D}_{\tilde{b}}\varphi_{1}|^{2}+|\varphi_{1}|^{4}+|\textrm{D}_{-a-\tilde{b}}\varphi_{2}|^{2}+|\varphi_{2}|^{4}-\frac{2\mathtt{i}}{4\pi}\tilde{b}\textrm{d}\tilde{b}-\frac{\mathtt{i}}{2\pi}\tilde{b}\textrm{d}a-\frac{\mathtt{i}}{4\pi}a\textrm{d}a.$ (111)
The phase transition is driven by the simultaneous condensation of $\varphi_{1}$ and $\varphi_{2}$.
In addition, several other dual critical theories based on level-rank dualities have been reviewed in Ref. [56].
Appendix F Static Structure Factor at Criticality
In this appendix, we provide technical details regarding the evaluation of integrals for the static structure factors, as expressed by Eq. (37) and Eq. (41), at the quantum critical points. If we naively consider the small-$\bm{k}$ expansion of the integrand before doing the integral, the leading-order term can integrated analytically, resulting in
$\Pi^{\tau\tau}(\tau\rightarrow 0,\bm{k})\approx\int\frac{\textrm{d}\omega}{2\pi}\frac{-|\bm{k}|^{2}k_{F}v_{F}(k_{F}v_{F}\sigma_{xx}^{b}+4\pi\omega(\sigma_{xx}^{b})^{2}-4\pi\omega(\sigma_{xy}^{b})^{2})}{\omega(k_{F}v_{F}+4\pi\omega\sigma_{xx}^{b}){}^{2}-16\pi^{2}\omega^{3}(\sigma_{xy}^{b})^{2}}=-\frac{\sigma_{xx}^{b}}{\pi}|\bm{k}|^{2}\log(1/|\bm{k}|)+\ldots.$ (112)
Here, a logarithmic divergence arises from the $\omega$-integral, leading to the factor $\log(1/|\bm{k}|)$ due to the identical scaling of $\omega$ and $\bm{k}$ at the critical point. The naive analytical calculation of Eq. (41) proceeds in a similar manner and is a special case of Eq. (112) with $\sigma_{xy}=0$.
Figure 4: (1) The numerical result (in black) of the static structure factor given by Eq. (37) at the CFL-FL Transition, where $k_{F}=\sqrt{2\pi}$, $v_{F}=0.25$, and $\sigma_{xx}=-\sigma_{xy}=1/(8\pi)$; (2) The numerical result (in black) of the static structure factor given by Eq. (41) at the Mott Transition, where $k_{F}=\sqrt{2\pi}$, $v_{F}=0.25$, and $\sigma_{xx}=0.355155/(2\pi)$. The numerical results are compared to the CFT-like scaling (in red) and that of ordinary Fermi surfaces (in blue).
To confirm the final result in Eq. (112), we also numerically evaluated the integrals using the exact expressions from Eq. (37) and Eq. (41), with the results presented in Fig. 4. In these numerical calculations, the value of $k_{F}$ is determined by the Luttinger theorem $V_{\textrm{FS}}/(2\pi)^{2}=1/2$ at half-filling [^6] The Luttinger theorem, as it applies to the composite Fermi liquid and the spinon Fermi surface, is discussed in Ref. [45]. The more precise expressions are $V_{\textrm{FS}}/(2\pi)^{2}=(1/2)/(2\pi\ell_{B}^{2})$ and $V_{\textrm{FS}}/(2\pi)^{2}=(1/2)/V_{\textrm{cell}}$, where $\ell_{B}$ represents the magnetic length and $V_{\textrm{cell}}$ denotes the volume of the unit cell. Our purpose is to numerically verify the analytic result in Eq. (112). To this end, we set $12\pi\ell_{B}^{2}=$ 1 and $V_{\textrm{cell}}=1$ for simplicity., where the Fermi-surface volume is given by $V_{\textrm{FS}}=\pi k_{F}^{2}$. We use a small value of $v_{F}$ for estimations. For the chargon conductivity $\sigma_{xx}^{b}$ at the Mott transition, we adopt the critical conductivity at the XY transition from conformal bootstrap [101]. For chargons at the CFL-FL transition, as a very crude estimation, we assume the universal transverse and Hall resistivities $\rho_{xx}^{b},\rho_{xy}^{b}$ are both around $4\pi$ [56]. In both cases, the numerical results agree well with the analytical expressions at small-$\bm{k}$, confirming the expectation that the static structure factor of critical Fermi surfaces behaves in a CFT-like manner and deviates from that of ordinary Fermi surfaces.
Appendix G Non-Linear Bosonization
In this appendix, we offer a brief introduction to non-linear bosonization [47] of Fermi liquids from the perspective of coherent-state construction. For any fermionic systems with translation symmetry, we can introduce the fermion bilinear operator
$\mathbbm{t}(\bm{x},\bm{k})=\int_{\bm{q}}c_{\bm{k}-\frac{\bm{q}}{2}}^{\dagger}c_{\bm{k}+\frac{\bm{q}}{2}}e^{\mathtt{i}\bm{q}\cdot\bm{x}},$ (113)
where $c$ is the gauge-invariant election operator. It generates an infinite-dimensional Lie algebra
$[\mathbbm{t}(\xi),\mathbbm{t}(\eta)]=-2\mathtt{i}\sin\left(\frac{\partial_{\xi}\curlywedge\partial_{\eta}}{2}\right)\delta^{2d}(\xi-\eta)\mathbbm{t}(\xi),$ (114)
[…]
where $|\textrm{FS}\rangle=\prod_{\epsilon_{\bm{k}}\leq\epsilon_{F}}c_{\bm{k}}^{\dagger}|0\rangle$ is the Fermi surface (FS) ground state. This is just the distribution function $F_{0}(\bm{x},\bm{k})$ that describes the shape of FS. This is analogous to magnetic orders that realize the condensation of fermion bilinears $\langle c^{\dagger}\vec{\sigma}c\rangle$ where $\vec{\sigma}$ are Pauli matrices.
The fluctuations of the condensation $\langle\mathbbm{t}\rangle$ are systematically described by nonlinear bosonization [47]. Given the symmetry group generated by the Lie algebra Eq. (114), one can introduce the coherent state
$|\phi\rangle=\exp\left(\mathtt{i}\int_{\bm{x},\bm{k}}\phi(\bm{x},\bm{k})\mathbbm{t}(\bm{x},\bm{k})\right)|\textrm{FS}\rangle,$ (116)
where $\phi(\bm{x},\bm{k})$ is a bosonic variable in phase space. The distribution function dressed by fluctuations is then
[…]
The quantization of FS fluctuations is given by the coherent-state path integral
$\mathcal{Z}=\int\mathcal{D}[\phi]e^{-\mathcal{S}[\phi]},\qquad\mathcal{S}[\phi]=\int\textrm{d}\tau\langle\phi|\partial_{\tau}+H|\phi\rangle.$ (118)
which can be unpacked order by order in terms of the boson $\phi$. In practical calculations [47], it is useful to consider a truncation of the algebra Eq. (114) using the separation of energy scales $\partial_{x}\partial_{k}\sim q/k_{F}\ll 1$, where $q$ is the low-energy relative momentum of particle-hole pairs. Another simplification comes from the redundancy in $\phi(\bm{x},\bm{k})$ due to the group generated by Eq. (114) being partially broken by the FS ground state. The FS fluctuations are sufficiently described by $\phi(\bm{x},\bm{k}_{F})$ where $\bm{k}_{F}$ labels points on the FS manifold [47]. The leading-order Gaussian part reproduces the well-known result based on patch assumptions
$\mathcal{S}=\frac{1}{4\pi}\int_{\bm{k}_{F}\in\textrm{FS}}\int_{\tau,\bm{x}}(\hat{\bm{v}}_{F}\cdot\partial_{\bm{x}}\phi)(\mathtt{i}\partial_{\tau}\phi+\bm{v}_{F}\cdot\partial_{\bm{x}}\phi),$ (119)
where $\bm{v}_{F}$ is the fermi velocity and $\hat{\bm{v}}_{F}$ denotes its direction. In other words, one has a chiral Luttinger liquid on each patch of the FS in the direction $\hat{\bm{v}}_{F}$. In the full action Eq. (118), different patches are allowed to talk to each other. Namely, the third-order term in $\phi$ contains the gradient $\partial_{\bm{k}_{F}}\phi$ along the FS, and therefore couples the nearest neighbor patches. For a more detailed analysis of the higher-order terms in the calculation, interested readers may consult Ref. [47]. It is worth noting that the procedure we have outlined here has parallels with the derivation of the non-linear sigma model for magnetic orders using spin coherent-state path integral methods. (see e.g. [102] for a textbook treatment).